### v1

# Newton polytopes in cluster algebras and $\tau$-tilting theory

Abstract.
We prove that the cluster monomials in non-initial cluster variables are uniquely determined by the Newton polytopes of their $F$-polynomials for skew-symmetrizable cluster algebras. Accordingly, we prove that the $\tau$-rigid modules and the left finite multi-simibricks in $\tau$-tilting theory are uniquely determined by their Newton polytopes of these modules.
2020 Mathematics Subject Classification:
13F60, 16G20
1. Introduction
Cluster algebras are a class of commutative algebras with an extra combinatorial structure introduced by Fomin and Zelevinsky [fz_2002]. Such algebras are generated by a special set of generators, called cluster variables, which are grouped into overlapping subsets of fixed size, called clusters. A seed is a pair consisting of a cluster ${\bf x}=(x_{1},\ldots,x_{n})$ and a skew-symmetrizable integer matrix $B$. New seeds can be obtained from a given one by a procedure called mutation. The sets of cluster variables and clusters of a cluster algebra are determined by an initial seed $({\bf x},B)$ and iterative mutations. A cluster monomial is a monomial in cluster variables from the same cluster. Cluster monomials are the central objects to study in cluster algebras.
The $g$-vectors and $F$-polynomials are introduced in [fomin_zelevinsky_2007] to study cluster monomials in terms of the initial seed $({\bf x},B)$. More precisely, each cluster monomial $u$ can be written as
$u={\bf x}^{{\bf g}_{u}}\cdot F_{u}(\hat{y}_{1},\ldots,\hat{y}_{n})\;\;\;\in\mathbb{Z}[x_{1}^{\pm 1},\ldots,x_{n}^{\pm 1}],$
where ${\bf g}_{u}\in\mathbb{Z}^{n}$ and $F_{u}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ are the $g$-vector and $F$-polynomial of $u$, and each $\hat{y}_{k}$ is a Laurent monomial in $x_{1},\ldots,x_{n}$. In particular, the $g$-vectors of the initial cluster variables $x_{1},\ldots,x_{n}$ are given by the columns of $I_{n}$ and the $F$-polynomials of $x_{1},\ldots,x_{n}$ are the constant $1$.
Fomin and Zelevinsky [fomin_zelevinsky_2007] conjectured that different cluster monomials have different $g$-vectors, which has been confirmed by Derksen, Weyman and Zelevinsky [DWZ10] for skew-symmetric cluster algebras and by Gross, Hacking, Keel, and Kontsevich for skew-symmetrizable cluster algebras.
$\tau$-tilting theory was introduced by Adachi, Iyama and Reiten [air_2014], which completes the classic tilting theory from the viewpoint of mutations. Various of the fundamental concepts in cluster algebras, such as clusters/seeds, mutations, $g$-vectors, $F$-polynomials etc., were generalized to $\tau$-tilting theory, cf. [air_2014, fei_2019a]. A dictionary of some concepts between cluster algebras and $\tau$-tilting theory are given in Table 1.
[…]
Table 1. Cluster algebras vs. $\tau$-tilting theory
It is known that $\tau$-rigid pairs in $\tau$-tilting theory play the role of cluster monomials in cluster algebras. Adachi, Iyama and Reiten [air_2014] proved that different $\tau$-rigid pairs have different $g$-vectors, which is analogous to the result that different cluster monomials have different $g$-vectors in cluster algebras.
Before we turn to the motivation part of this paper. Let us give the definition of Newton polytopes of polynomials and modules used in this paper.
Definition 1.1 (Newton polytopes of polynomials and modules).
(i) Let $F({\bf y})=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}{\bf y}^{\bf v}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ be a non-zero polynomial. Its Newton polytope $\mathsf{P}(F)$ is defined to be the convex hull of the finite set $\{{\bf v}\in\mathbb{N}^{n}\mid c_{\bf v}\neq 0\}$.
(ii) The Newton polytope $\mathsf{P}(M)$ of a module $M\in\operatorname{\mathsf{mod}}\nolimits A$ is defined to be the convex hull of the dimension vectors of the quotient modules of $M$.
Recently, there has been growing interest in the study of $F$-polynomials. For example, Jiarui Fei [fei_2019a, fei_2019b] studied the tropical $F$-polynomials of modules. In particular, he proved that the Newton polytope of the $F$-polynomial $F_{M}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ of a module $M\in\operatorname{\mathsf{mod}}\nolimits A$ coincides with the Newton polytope of $M$. The $F$-invariant in cluster algebras is introduced in [Cao-2023] as a generalization of the $E$-invariant [DWZ10] in additive cluster categorification and the $\mathfrak{d}$-invariant [kkko-2018, kkop-2020] in monoidal cluster categorification. It turns out that the $F$-invariant is related to the tropical $F$-polynomials.
Since the $g$-vectors can determine the cluster monomials and $\tau$-rigid pairs, it is natural to ask whether the $F$-polynomials can determine the cluster monomials and $\tau$-rigid pairs?
In order to consider this question, it is natural to exclude the initial cluster variables $x_{1},\ldots,x_{n}$ and correspondingly the initial indecomposable $\tau$-rigid pairs $(0,P_{1}),\ldots,(0,P_{n})$, because the $F$-polynomials of these initial objects are $1$.
In [ckq_2024], Keller, Qin and the author of the present paper proved that cluster monomials in non-initial cluster variables are uniquely determined by their $F$-polynomials. The proof there depends on a notion of valuation pairing on (upper) cluster algebras. Typically, the same proof does not work in $\tau$-tilting theory.
Recently, it is proved [cao-2025b] that the indecomposable $\tau$-rigid modules and left finite bricks in $\operatorname{\mathsf{mod}}\nolimits A$ are uniquely determined by their Newton polytopes, which is equivalent to saying that such modules are uniquely determined by the Newton polytopes of their $F$-polynomials, thanks to Fei’s result [fei_2019a]. The proof uses two key results: (i) all functorially finite torsion classes are semistable torsion classes [asai-2021]; and (ii) the brick-$\tau$-rigid correspondence in [DIJ-17], which is a bijection between the indecomposable $\tau$-rigid modules and the left finite bricks. In the general case, it is proved that if two rigid modules $U$ and $V$ (not necessarily indecomposable) have the same Newton polytope, then $U\oplus V$ remains $\tau$-rigid. Unfortunately, it is not clear from [cao-2025b] whether $U$ and $V$ are isomorphic or not. The main motivation for this paper is the following question.
Question 1.2.
Whether the Newton polytopes of $F$-polynomials determine the cluster monomials in non-initial cluster variables and $\tau$-rigid modules?
Recently, the Newton polytopes of $F$-polynomials and modules were extensively studied in the literature, cf. [BK-2012, BKT-2014, BCDMTY-2024, fei_2019a, fei_2019b, AHIKM-2022, Asai-Iyama-2020, li-pan-2022, cao-2025b]. In particular, the Newton polytopes of modules over preprojective algebras are used to study MV polytopes in [BKT-2014], which parametrize Lusztig’s canonical basis, cf. [Kamnitzer-2010].
Before giving the main results of this paper, we first recall the definitions of $\tau$-rigid modules and left finite multi-semibricks.
Definition 1.3 ($\tau$-rigid module, multi-semibrick and left finite module).
(i) A module $M\in\operatorname{\mathsf{mod}}\nolimits A$ is called $\tau$-rigid, if $\operatorname{\mathsf{Hom}}_{A}(M,\tau M)=0$, where $\tau$ is the Auslander-Reiten translation in $\operatorname{\mathsf{mod}}\nolimits A$.
(ii) A module $M\in\operatorname{\mathsf{mod}}\nolimits A$ is called a multi-semibrick if $M$ has a decomposition $M\cong\oplus_{i=1}^{r}C_{i}^{a_{i}}$ such that each $C_{i}\in\operatorname{\mathsf{mod}}\nolimits A$ is a brick and $\operatorname{\mathsf{Hom}}_{A}(C_{i},C_{j})=0$ for any $i\neq j$.
(iii) A module $M\in\operatorname{\mathsf{mod}}\nolimits A$ is left finite, if the smallest torsion class containing $M$ is a functorially finite torsion class in $\operatorname{\mathsf{mod}}\nolimits A$.
The main results in this paper are as follows.
Theorem 1.4.
Let $\mathcal{A}$ be a skew-symmetrizable cluster algebra with initial seed $({\bf x}_{t_{0}},B_{t_{0}})$. Let $u$ and $v$ be two cluster monomials in non-initial cluster variables. If the two $F$-polynomials $F_{u}^{t_{0}}$ and $F_{v}^{t_{0}}$ have the same Newton polytope, then $u=v$.
Theorem 1.5.
Let $A$ be a finite dimensional basic algebra over an algebraically closed field $\mathbf{k}$. Let $U$ and $V$ be two $\tau$-rigid modules in $\operatorname{\mathsf{mod}}\nolimits A$. If $U$ and $V$ have the same Newton polytope, then $U\cong V$.
The left finite multi-semibricks are the dual side of $\tau$-rigid modules in some sense. So it is natural to extend the result on $\tau$-rigid modules to the left finite multi-semibricks.
Theorem 1.6.
Let $U$ and $V$ be two left finite multi-semibricks in $\operatorname{\mathsf{mod}}\nolimits A$. If $U$ and $V$ have the same Newton polytope, then $U\cong V$.
The proofs of Theorem 1.4 and Theorem 1.5 are quite similar, whereas the proof of Theorem 1.6 is slightly different. Let make some comments about the proofs of Theorem 1.4 and Theorem 1.5. Both proofs are based on the reduction arguments. The key points are summarized as follows:
  • (i)
    Show that the direct sum $U\oplus V$ remains a $\tau$-rigid module; and the product $uv$ remains a cluster monomial.
  • (ii)
    The proof of Theorem 1.5 is by reducing the number $|U\oplus V|$ of iso-classes of indecomposable direct summands of $U\oplus V$. The proof of Theorem 1.4 is by reducing the number $|\operatorname{\mathsf{supp}}(uv)|$ of the support set $\operatorname{\mathsf{supp}}(uv)$ of $uv$, which is the set of cluster variables appearing in the cluster monomial uvuv.
  • (iii)
    In order to reduce the number $|U\oplus V|$, we consider the left Bongartz completion $\mathcal{M}=(M,P)$ of $U\oplus V$, which is the basic $\tau$-tilting pair such that $\operatorname{\mathsf{Fac}}M=\operatorname{\mathsf{Fac}}(U\oplus V)$. Then we show that any left mutation of $\mathcal{M}=(M,P)$ can be used to construct two new $\tau$-rigid modules $U^{\prime}\in\operatorname{\mathsf{add}}\nolimits U$ and $V^{\prime}\in\operatorname{\mathsf{add}}\nolimits V$ satisfying that $U^{\prime}$ and $V^{\prime}$ have the same Newton polytope but $|U^{\prime}\oplus V^{\prime}|<|U\oplus V|$. Moreover, $U\cong V$ if and only if $U^{\prime}\cong V^{\prime}$.
  • (iv)
    In order to reduce the number $|\operatorname{\mathsf{supp}}(uv)|$, we consider the left Bongartz completion ${\bf x}_{s}]$ of the partial cluster $\operatorname{\mathsf{supp}}(uv)$. Then we show that any red mutation of the seed $({\bf x}_{s},B_{s})$ can be used to construct two new cluster monomials $u^{\prime}$ and $v^{\prime}$ satisfying that $u^{\prime}$ and $v^{\prime}$ have the same Newton polytope but $\operatorname{\mathsf{supp}}(u^{\prime}v^{\prime})\subsetneq\operatorname{\mathsf{supp}}(uv)$. Moreover, $u=v$ if and only if $u^{\prime}=v^{\prime}$.
Of course, many preparations are required to ensure that each stage goes smoothly. These will be detailed in the main body of the paper. Among these, the $F$-invariant and partial $F$-invariant in cluster algebra and $\tau$-tilting theory will play a crucial role in the proofs, because such invariants are directly related to the Newton polytopes. We give a detailed discussion on $F$-invariant of cluster monomials in Subsection 4.1 and that of decorated modules in Subsection 5.1.
This paper contains two closely related parts, cluster algebras and $\tau$-tilting theory. These two parts can be read in any order. Personally, I first proved the main result in cluster algebras and then realized that a similar approach can be applied to the $\tau$-tilting theory.
Remark 1.7.
The prefixes “left” and “right” in the terminology “left mutation”, “right mutation”, “left Bongartz completion”, and “right Bongartz completion” indicate the direction of change under a suitable partial order: “left” denotes a move to a smaller element, while “right” denotes a move to a larger one. This convention aligns with the familiar ordering of integers $(\dots,-2,-1,0,1,2,\ldots)$, where moving left yields smaller numbers and moving right yields larger ones.
Acknowledgement. The author is partially supported the National Key R&D Program of China (2024YFA1013801).
2. Preliminaries on polytopes and cluster algebras
Throughout, we denote by $\langle{-,-}\rangle:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ the standard inner product on $\mathbb{R}^{n}$ and by ${\bf e}_{1},\ldots,{\bf e}_{n}$ the standard basis of $\mathbb{R}^{n}$.
2.1. Polytopes, Minkowski sum and tropical polynomials
A polytope $\mathsf{P}$ in $\mathbb{R}^{n}$ is the convex hull of a finite (non-empty) subset of $\mathbb{R}^{n}$, which is a bounded closed subset in $\mathbb{R}^{n}$. For a polytope $\mathsf{P}$ in $\mathbb{R}^{n}$, its support function $h_{\mathsf{P}}:\mathbb{R}^{n}\to\mathbb{R}$ is defined by
$h_{\mathsf{P}}({\bf r}):=\max\{\langle{{\bf a},{\bf r}}\rangle\mid{\bf a}\in\mathsf{P}\}.$
It is known from [BK-2012, Section 4.2] or [Schneider_2013, Section 1.7] that the polytope $\mathsf{P}$ can be recovered from its support function $h_{\mathsf{P}}$ by
$\mathsf{P}=\{{\bf a}\in\mathbb{R}^{n}\mid\langle{{\bf a},{\bf r}}\rangle\leq h_{\mathsf{P}}({\bf r}),\;\forall{\bf r}\in\mathbb{R}^{n}\}.$
Let $\mathsf{P}_{1}$ and $\mathsf{P}_{2}$ be two polytopes in $\mathbb{R}^{n}$. The Minkowski sum of $\mathsf{P}_{1}$ and $\mathsf{P}_{2}$ is the polytope in $\mathbb{R}^{n}$ given by
$\mathsf{P}_{1}+\mathsf{P}_{2}:=\{{\bf a}+{\bf b}\mid{\bf a}\in\mathsf{P}_{1},\;{\bf b}\in\mathsf{P}_{2}\}.$
Theorem 2.1 ([Schneider_2013, Theorem 1.7.5]).
Let $\mathsf{P}_{1}$ and $\mathsf{P}_{2}$ be two polytopes in $\mathbb{R}^{n}$. Then
$h_{\mathsf{P}_{1}+\mathsf{P}_{2}}=h_{\mathsf{P}_{1}}+h_{\mathsf{P}_{2}}.$
As a direct consequence, the following cancellation law holds.
Corollary 2.2.
Let $\mathsf{P}_{1},\mathsf{P}_{2}$ and $\mathsf{Q}$ be three polytopes in $\mathbb{R}^{n}$. If $\mathsf{P}_{1}+\mathsf{Q}=\mathsf{P}_{2}+\mathsf{Q}$, then $\mathsf{P}_{1}=\mathsf{P}_{2}$.
Proof.
Since $\mathsf{P}_{1}+\mathsf{Q}=\mathsf{P}_{2}+\mathsf{Q}$ and by Theorem 2.1, we have
$h_{\mathsf{P}_{1}}+h_{\mathsf{Q}}=h_{\mathsf{P}_{1}+\mathsf{Q}}=h_{\mathsf{P}_{2}+\mathsf{Q}}=h_{\mathsf{P}_{2}}+h_{\mathsf{Q}}.$
Thus $h_{\mathsf{P}_{1}}=h_{\mathsf{P}_{2}}$. Since a polytope is uniquely determined its support function, we get $\mathsf{P}_{1}=\mathsf{P}_{2}$. ∎
In this paper, we mainly focus on the polytopes defined from polynomials and modules of a finite dimension algebra $A$.
Definition 2.3 (Newton polytopes of polynomials and modules).
(i) Let $F({\bf y})=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}{\bf y}^{\bf v}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ be a non-zero polynomial. Its Newton polytope $\mathsf{P}(F)$ is defined to be the convex hull of the finite set $\{{\bf v}\in\mathbb{N}^{n}\mid c_{\bf v}\neq 0\}$.
(ii) The Newton polytope $\mathsf{P}(M)$ of a module $M\in\operatorname{\mathsf{mod}}\nolimits A$ is defined to be the convex hull of the dimension vectors of the quotient modules of $M$.
Proposition 2.4 ([GKZ-1994, Chapter 6, Prop. 1.2]).
Let $F_{1}$ and $F_{2}$ be two non-zero polynomials in $\mathbb{Z}[y_{1},\ldots,y_{n}]$. Then
$\mathsf{P}(F_{1}F_{2})=\mathsf{P}(F_{1})+\mathsf{P}(F_{2}),$
where $\mathsf{P}(F_{k})$ is the Newton polytope of $F_{k}$ for $k=1,2$.
Definition 2.5 (Tropical polynomial).
Let $F({\bf y})=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}{\bf y}^{\bf v}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ be a non-zero polynomial. The tropical polynomial of $F$ is the map $F[-]:\mathbb{R}^{n}\to\mathbb{R}$ defined by
$F[{\bf r}]:=\max\{\langle{{\bf v},{\bf r}}\rangle\mid c_{\bf v}\neq 0\}.$
Tropical polynomials play an important role in defining the $F$-invariant in cluster algebras and $\tau$-tilting theory in Subsections 4.1, 5.1. We can see that if ${\bf r}\in\mathbb{Z}^{n}$, then $F[{\bf r}]\in\mathbb{Z}$.
Remark 2.6.
Note that if the polynomial $F$ has constant term $1$, then $F[{\bf r}]\in\mathbb{Z}_{\geq 0}$ for any ${\bf r}\in\mathbb{Z}^{n}$. We remark that the $F$-polynomials of cluster monomials and modules always have constant term $1$.
Example 2.7.
Take $F=1+y_{1}+y_{1}y_{2}\in\mathbb{Z}[y_{1},y_{2}]$ and ${\bf r}=\begin{bmatrix}-2\\ 1\end{bmatrix}$, then
$F[{\bf r}]=\max\{\begin{bmatrix}0,0\end{bmatrix}\begin{bmatrix}-2\\ 1\end{bmatrix},\begin{bmatrix}1,0\end{bmatrix}\begin{bmatrix}-2\\ 1\end{bmatrix},\begin{bmatrix}1,1\end{bmatrix}\begin{bmatrix}-2\\ 1\end{bmatrix}\}=\max\{0,-2,-1\}=0.$
The following result can be checked easily.
Proposition 2.8.
Let $F\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ be a non-zero polynomial and $\mathsf{P}(F)$ its Newton polytope. Then the tropical polynomial of $F$ and the support function of $\mathsf{P}(F)$ are the same, i.e.,
$F[{\bf r}]=h_{\mathsf{P}(F)}({\bf r}),\quad\forall{\bf r}\in\mathbb{R}^{n}.$
Corollary 2.9.
Let $F_{1},F_{2}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ be two non-zero polynomials. Then for any ${\bf r}\in\mathbb{R}^{n}$, we have $(F_{1}F_{2})[{\bf r}]=F_{1}[{\bf r}]+F_{2}[{\bf r}]$.
Proof.
By Proposition 2.4, we know that $\mathsf{P}(F_{1}F_{2})=\mathsf{P}(F_{1})+\mathsf{P}(F_{2})$. Then by Theorem 2.1, we have
$h_{\mathsf{P}(F_{1}F_{2})}=h_{\mathsf{P}(F_{1})+\mathsf{P}(F_{2})}=h_{\mathsf{P}(F_{1})}+h_{\mathsf{P}(F_{2})}.$
Then the desired result follows from Proposition 2.8. ∎
2.2. Cluster algebras
We first recall Fomin-Zelevinsky’s matrix mutation [fz_2002].
Definition 2.10 (Matrix mutation).
Let $A=(a_{ij})$ be an $m\times n$ integer matrix. For any integer $k$ with $k\leq m$ and $k\leq n$, the mutation of $A$ in direction $k$ is defined to be the new integer matrix $\mu_{k}(A)=A^{\prime}=(a_{ij}^{\prime})$ given by
$a_{ij}^{\prime}$ $=$ $\begin{cases}-a_{ij},&\text{if}\;i=k\;\text{or}\;j=k,\\ a_{ij}+[a_{ik}]_{+}[a_{kj}]_{+}-[-a_{ik}]_{+}[-a_{kj}]_{+},&\text{otherwise},\end{cases}$
where $a]_{+}:=\max\{a,0\}$ for any $a\in\mathbb{R}$.
Now we fix a positive integer $n$ and denote by $1,n]:=\{1,2,\ldots,n\}$. An $n\times n$ integer matrix $B$ is said to be skew-symmetrizable, if there exists a diagonal integer matrix $D=diag(d_{1},\ldots,d_{n})$ with each $d_{i}>0$ such that $DB$ is skew-symmetric. Such a diagonal matrix $D$ is called a skew-symmetrizer of $B$.
Proposition 2.11 ([fz_2002]).
(i) For any integer matrix $A=(a_{ij})_{m\times n}$, we have $\mu_{k}^{2}(A)=A$.
(ii) If $B=(b_{ij})_{n\times n}$ is skew-symmetrizable, then $B^{\prime}:=\mu_{k}(B)$ is still skew-symmetrizable and the two matrices $B,B^{\prime}$ share the same skew-symmetrizers.
A seed in $\mathbb{F}:=\mathbb{Q}(z_{1},\ldots,z_{n})$ is a pair $({\bf x},B)$, where
  • ${\bf x}=(x_{1},\ldots,x_{n})$ is an ordered set of free generators of $\mathbb{F}$ over $\mathbb{Q}$;
  • $B=(b_{ij})$ is an $n\times n$ skew-symmetrizable matrix.
Definition 2.12 (Seed mutation).
Let $({\bf x},B)$ be a seed in $\mathbb{F}$. The mutation of $({\bf x},B)$ in direction $k\in[1,n]$ is the new seed $({\bf x}^{\prime},B^{\prime})=\mu_{k}({\bf x},B)$ given by $B^{\prime}=\mu_{k}(B)$ and
$x_{i}^{\prime}=\begin{cases}x_{i},&i\neq k,\\ x_{k}^{-1}\cdot(\prod_{j=1}^{n}x_{j}^{[b_{jk}]_{+}}+\prod_{j=1}^{n}x_{j}^{[-b_{jk}]_{+}}),&i=k.\end{cases}$
It can be checked that $\mu_{k}$ is an involution. Let $\mathbb{T}_{n}$ denote the $n$-regular tree. We label the edges of $\mathbb{T}_{n}$ by $1,\ldots,n$ such that the $n$ different edges adjacent to the same vertex of $\mathbb{T}_{n}$ receive different labels.
Definition 2.13 (Cluster pattern).
A cluster pattern $\mathcal{S}_{X}=\{({\bf x}_{t},B_{t})\mid t\in\mathbb{T}_{n}\}$ is an assignment of a seed $({\bf x}_{t},B_{t})$ in $\mathbb{F}$ to every vertex $t$ of $\mathbb{T}_{n}$ such that $({\bf x}_{t^{\prime}},B_{t^{\prime}})=\mu_{k}({\bf x}_{t},B_{t})$ whenever ${t}$${t^{\prime}}$$\scriptstyle{k}$ in $\mathbb{T}_{n}$.
Usually, we fix a vertex $t_{0}\in\mathbb{T}_{n}$ as the rooted vertex of $\mathbb{T}_{n}$. The seed of a cluster pattern at the rooted vertex $t_{0}$ is called an initial seed. We call ${\bf x}_{t}$ and $B_{t}$ the cluster and exchange matrix at the vertex $t\in\mathbb{T}_{n}$ and write ${\bf x}_{t}=(x_{1;t},\ldots,x_{n;t})$ and $B_{t}=(b_{ij}^{t})$. Elements in clusters are called cluster variables.
The cluster algebra $\mathcal{A}$ associated to a cluster pattern $\mathcal{S}_{X}=\{({\bf x}_{t},B_{t})\mid t\in\mathbb{T}_{n}\}$ is the $\mathbb{Z}$-subalgebra of $\mathbb{F}=\mathbb{Q}(z_{1},\ldots,z_{n})$ given by
$\mathcal{A}=\mathbb{Z}[x_{1;t},\ldots,x_{n;t}\mid t\in\mathbb{T}_{n}].$
A cluster monomial $u$ of $\mathcal{A}$ is a monomial in cluster variables from the same cluster, i.e.,
$u={\bf x}_{t}^{\bf h}=x_{1;t}^{h_{1}}\cdots x_{n;t}^{h_{n}}$
for some vertex $t\in\mathbb{T}_{n}$ and ${\bf h}=(h_{1},\ldots,h_{n})^{T}\in\mathbb{N}^{n}$.
Theorem 2.14 (Laurent phenomenon and separation formula [fz_2002, fomin_zelevinsky_2007]).
Let $\mathcal{A}$ be a cluster algebra with initial seed $({\bf x}_{t_{0}},B_{t_{0}})$. Then the following statements hold.
  • (i)
    Any cluster monomial $u$ can be written as a Laurent polynomial in $\mathbb{Z}[x_{1;t_{0}}^{\pm 1},\ldots,x_{n;t_{0}}^{\pm 1}]$.
  • (ii)
    Denote by $\hat{y}_{k;t_{0}}={\bf x}_{t_{0}}^{B_{t_{0}}{\bf e}_{k}}$. The Laurent polynomial in (i) has a canonical expression
    $u={\bf x}_{t_{0}}^{{\bf g}_{u}^{t_{0}}}F_{u}^{t_{0}}(\hat{y}_{1;t_{0}},\ldots,\hat{y}_{n;t_{0}})\;\;\;\in\;\;\mathbb{Z}[x_{1;t_{0}}^{\pm 1},\ldots,x_{n;t_{0}}^{\pm 1}],$
    where $g_{u}^{t_{0}}\in\mathbb{Z}^{n}$ and $F_{u}^{t_{0}}({\bf y})\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ are canonically defined from principal cluster algebras [fomin_zelevinsky_2007, (6.4) & (3.3)].
Definition 2.15 ($g$-vector, $G$-matrix and $F$-polynomial).
Let $u$ be a cluster monomial of $\mathcal{A}$ and keep the notations above.
  • (i)
    The integer vector ${\bf g}_{u}^{t_{0}}\in\mathbb{Z}^{n}$ is called the $g$-vector of $u$ with respect to (the seed at) vertex $t_{0}$.
  • (ii)
    The matrix $G_{t}^{t_{0}}=({\bf g}_{x_{1;t}}^{t_{0}},\ldots,{\bf g}_{x_{n;t}}^{t_{0}})$ is called the $G$-matrix of $({\bf x}_{t},B_{t})$ with respect to vertex $t_{0}$.
  • (iii)
    The polynomial $F_{u}^{t_{0}}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ is called the $F$-polynomial of $u$ with respect to vertex $t_{0}$.
Example 2.16 (Cluster algebra of type $A_{2}$).
Take $B=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$ and ${\bf x}=(x_{1},x_{2})$. It is easy to check that the cluster algebra $\mathcal{A}$ defined by the initial seed $({\bf x},B)$ has only five (unlabeled) clusters
$\{x_{1},x_{2}\},\;\{x_{2},x_{3}\},\;\{x_{3},x_{4}\},\;\{x_{4},x_{5}\},\;\{x_{5},x_{1}\},\quad\text{where}$
$x_{3}:=\frac{x_{2}+1}{x_{1}},\;x_{4}:=\frac{x_{1}+x_{2}+1}{x_{1}x_{2}},\;x_{5}:=\frac{x_{1}+1}{x_{2}}.$
The canonical expressions of the three non-initial cluster variables with respect to the initial seed $({\bf x},B)$ are given as follows:
$x_{3}={x_{1}^{-1}x_{2}\cdot(1+\widehat{y}_{1}),}\;\;\;x_{4}={x_{1}^{-1}\cdot(1+\widehat{y}_{1}+\widehat{y}_{1}\widehat{y}_{2}),}\;\;\;x_{5}={x_{2}^{-1}\cdot(1+\widehat{y}_{2})},$
where ${\hat{y}}_{1}={\bf x}^{B{\bf e}_{1}}=x_{2}^{-1}$ and ${\hat{y}}_{2}={\bf x}^{B{\bf e}_{2}}=x_{1}$.
Let $\mathcal{A}$ be a cluster algebra whose seed at vertex $t\in\mathbb{T}_{n}$ is denoted by $({\bf x}_{t},B_{t})$. Let $t_{0}$ and $t$ be two vertices of $\mathbb{T}_{n}$ and $\overleftarrow{\mu}$ the mutation sequence corresponding to the unique path from the vertex $t_{0}$ to $t$ in $\mathbb{T}_{n}$. We apply the mutation sequence $\overleftarrow{\mu}$ to $\begin{pmatrix}B_{t_{0}}\\ I_{n}\end{pmatrix}$, then the resulting matrix $\overleftarrow{\mu}\begin{pmatrix}B_{t_{0}}\\ I_{n}\end{pmatrix}$ takes the form $\begin{pmatrix}B_{t}\\ C_{t}^{t_{0}}\end{pmatrix}$ for some $n\times n$ integer matrix $C_{t}^{t_{0}}$.
Definition 2.17 ($C$-matrix and $c$-vector).
Keep the above notations. We call the $n\times n$ integer matrix $C_{t}^{t_{0}}$ the $C$-matrix of $({\bf x}_{t},B_{t})$ with respect to vertex $t_{0}\in\mathbb{T}_{n}$, whose columns are called $c$-vectors.
Theorem 2.18 ([GHKK18],[NZ12]).
Let $\mathcal{A}$ be a cluster algebra with initial seed $({\bf x}_{t_{0}},B_{t_{0}})$. The following statements hold.
  • (i)
    For any vertex $t\in\mathbb{T}_{n}$, we have $(G_{t}^{t_{0}})^{T}DC_{t}^{t_{0}}D^{-1}=I_{n}$, where $D$ is a skew-symmetrizer for the exchange matrices of $\mathcal{A}$.
  • (ii)
    Each row vector of a $G$-matrix $G_{t}^{t_{0}}$ is either non-negative or non-positive.
  • (iii)
    Each column vector of a $C$-matrix $C_{t}^{t_{0}}$ is either non-negative or non-positive.
  • (iv)
    The $F$-polynomial $F_{u}^{t_{0}}({\bf y})$ of a cluster monomial $u$ is a polynomial in $\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$ with constant term $1$.
Definition 2.19 (Green and red mutation).
Let $\mathcal{A}$ be a cluster algebra with initial seed $({\bf x}_{t_{0}},B_{t_{0}})$. A seed mutation $\mu_{k}({\bf x}_{t},B_{t})$ in $\mathcal{A}$ is called a green mutation, if the $k$-th column of the $C$-matrix $C_{t}=C_{t}^{t_{0}}$ is a non-negative vector. Otherwise, it is called a red mutation.
Denote by ${\bf x}_{t}]$ the cluster ${\bf x}_{t}=(x_{1;t},\cdots,x_{n;t})$ up to permutations, that is, $\mathbf{x}_{t}]=\{x_{1;t},\cdots,x_{n;t}\}$.
Proposition 2.20.
Let $\mathcal{A}$ be a cluster algebra initial seed $({\bf x}_{t_{0}},B_{t_{0}})$. If the $C$-matrix $C_{t}^{t_{0}}$ is a non-negative matrix, then ${\bf x}_{t}]=[{\bf x}_{t_{0}}]$.
Proof.
Let $\check{\mathcal{A}}$ be the cluster algebra with initial seed $({\bf z}_{t_{0}},B_{t_{0}}^{T})$ and we write $({\bf x}_{t},B_{t}^{T})$ for its seed at vertex $t$. We use $\check{G}_{t_{2}}^{t_{1}}$ to denote the $G$-matrix of $({\bf z}_{t_{2}},B_{t_{2}}^{T})$ with respect to vertex $t_{1}$. By [NZ12, (1.13)], we have $C_{t}^{t_{0}}=(\check{G}_{t_{0}}^{t})^{T}$. Since $C_{t}^{t_{0}}$ is a non-negative matrix, we know that $\check{G}_{t_{0}}^{t}$ is a non-negative matrix. Now we take $({\bf z}_{t},B_{t}^{T})$ as the initial seed of $\check{\mathcal{A}}$. The non-negative $G$-matrix $\check{G}_{t_{0}}^{t}$ corresponds to the positive chamber in the scattering diagram [GHKK18] of $\check{\mathcal{A}}$. This implies ${\bf z}_{t_{0}}]=[{\bf z}_{t}]$. So $\check{G}_{t_{0}}^{t}$ is a permutation matrix. Thus $C_{t}^{t_{0}}$ is also a permutation matrix. Therefore, $(C_{t}^{t_{0}})^{-1}$ is a permutation matrix. In particular, it is a non-negative matrix. Then by Theorem 2.18 (i), we see that the $G$-matrix $G_{t}^{t_{0}}=D(C_{t}^{t_{0}})^{-1}D^{-1}$ of $\mathcal{A}$ is a non-negative matrix. So it corresponds to the positive chamber in the scattering diagram of $\mathcal{A}$. Thus ${\bf x}_{t}]=[{\bf x}_{t_{0}}]$. ∎
2.3. Bongartz completion in cluster algebras
A partial cluster of $\mathcal{A}$ is a subset of some cluster of $\mathcal{A}$. Recall that we denote by $\mathbf{x}_{t}]=\{x_{1;t},\cdots,x_{n;t}\}$.
Definition 2.21 (Left and right Bongartz completion, [CGY-2023]).
Let $\mathcal{A}$ be a cluster algebra and $U$ a partial cluster of $\mathcal{A}$.
  • (i)
    A cluster ${\bf x}_{s}]$ is called the left Bongartz completion of $U$ with respect to a vertex $t_{0}\in\mathbb{T}_{n}$ if the following two conditions hold.
    • (a)
      $U$ is a subset of $\mathbf{x}_{s}]$;
    • (b)
      The $i$-th column of the $C$-matrix $C_{s}^{t_{0}}$ is a non-negative vector for any $i$ such that $x_{i;s}\notin U$.
  • (ii)
    A cluster ${\bf x}_{s}]$ is called the right Bongartz completion of $U$ with respect to vertex $t_{0}\in\mathbb{T}_{n}$ if the following two conditions hold.
    • (a)
      $U$ is a subset of $\mathbf{x}_{s}]$;
    • (b)
      The $i$-th column of the $C$-matrix $C_{s}^{t_{0}}$ is a non-positive vector for any $i$ such that $x_{i;s}\notin U$.
Remark 2.22.
The left Bongartz completion and right Bongartz completion are called the Bongartz completion and Bongartz co-completion in [CGY-2023].
Notice that both the existence and uniqueness of left and right Bongartz completion in cluster algebras are not clear from their own definitions. Let us look at the case $U=\emptyset$. In this case, we have the following facts:
  • A cluster ${\bf x}_{s}]$ is the left Bongartz completion of $U=\emptyset$ with respect to vertex $t_{0}$ if and only if the $C$-matrix $C_{s}^{t_{0}}$ is a non-negative matrix, which implies that ${\bf x}_{s}]=[{\bf x}_{t_{0}}]$, by Proposition 2.20.
  • A cluster ${\bf x}_{s}]$ is the right Bongartz completion of $U=\emptyset$ with respect to vertex $t_{0}$ if and only if the $C$-matrix $C_{s}^{t_{0}}$ is a non-positive matrix. Notice that such a $C$-matrix exists if and only if the exchange matrix $B_{t_{0}}$ has a green-to-red sequence in the sense of [Muller16]. In particular, this implies that the right Bongartz completion might not exist in general.
Theorem 2.23 ([CGY-2023, Theorem 4.15]).
Let $\mathcal{A}$ be a cluster algebra with initial seed $({\bf x}_{t_{0}},B_{t_{0}})$. Then for any partial cluster $U$ of $\mathcal{A}$, there exists a unique cluster ${\bf x}_{s}]$ such that ${\bf x}_{s}]$ is the left Bongartz completion of $U$ with respect to vertex $t_{0}$.
Remark 2.24.
Note that the right Bongartz completion in cluster algebras might not exist in general. But if it exists, it is unique, by [CGY-2023, Corollary 5.5]. For this paper, we only need to use the left Bongartz completion, which always exists by Theorem 2.23.
Recall that for a nonzero polynomial $F({\bf y})=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}{\bf y}^{\bf v}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$, its tropical polynomial $F[-]:\mathbb{R}^{n}\to\mathbb{R}$ is defined by
$F[{\bf r}]:=\max\{\langle{{\bf v},{\bf r}}\rangle\mid c_{\bf v}\neq 0\}.$
Denote by $\mathcal{X}$ the set of cluster variables of $\mathcal{A}$. Given a cluster monomial $u={\bf x}_{t}^{\bf h}=\prod x_{i;t}^{h_{i}}$ in seed $({\bf x}_{t},B_{t})$, we define its dominant set with respect to the initial seed $({\bf x}_{t_{0}},B_{t_{0}})$ as follows:
$\operatorname{\mathsf{dom}}^{t_{0}}(u)=\{z\in\mathcal{X}\mid F_{z}^{t_{0}}[D{\bf g}_{u}^{t_{0}}]=0\}.$
Since the $F$-polynomials of initial cluster variables are $1$, we see that the initial cluster variables are always contained in $\operatorname{\mathsf{dom}}^{t_{0}}(u)$.
Definition 2.25 (Dominant set of a seed).
Let $({\bf x}_{t},B_{t})$ be a seed of $\mathcal{A}$. The dominant set $\operatorname{\mathsf{dom}}^{t_{0}}[t]$ of $({\bf x}_{t},B_{t})$ with respect to the initial seed $({\bf x}_{t_{0}},B_{t_{0}})$ is defined to be the dominant set of the multiplicity free cluster monomial $u_{t}=\prod_{i=1}^{n}x_{i;t}$ with full support, that is,
$\operatorname{\mathsf{dom}}^{t_{0}}[t]\coloneqq\operatorname{\mathsf{dom}}^{t_{0}}(u_{t})=\{z\in\mathcal{X}\mid F_{z}^{t_{0}}[D{\bf g}_{u_{t}}^{t_{0}}]=0\}.$
Remark 2.26.
The dominant sets are introduced in [Cao-2023] as a replacement of torsion classes in the $\tau$-tilting theory. The non-initial cluster variables in the dominant set $\operatorname{\mathsf{dom}}^{t_{0}}[t]$ correspond to the indecomposable $\tau$-rigid modules contained in the torsion class $\operatorname{\mathsf{Fac}}M=\prescript{\bot}{}{(}{\tau M})\cap P^{\bot}$ for a $\tau$-tilting pair $(M,P)$ in the $\tau$-tilting theory, cf. [Cao-2023, Proposition 7.23] or Corollary 5.11 in this paper.
Theorem 2.27 ([cao-2026a]).
Let $\mathcal{A}$ be a cluster algebra with initial seed $({\bf x}_{t_{0}},B_{t_{0}})$ and $U$ a partial cluster of $\mathcal{A}$. Then the following statements hold.
  • (i)
    A cluster ${\bf x}_{s}]$ is the left Bongartz completion of $U$ with respect to vertex $t_{0}$ if and only if $U\subseteq[{\bf x}_{s}]$ and $\operatorname{\mathsf{dom}}^{t_{0}}[s]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t]$ for any cluster ${\bf x}_{t}]$ with $U\subseteq[{\bf x}_{t}]$.
  • (ii)
    A cluster ${\bf x}_{s}]$ is the right Bongartz completion of $U$ with respect to vertex $t_{0}$ if and only if $U\subseteq[{\bf x}_{s}]$ and $\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[s]$ for any cluster ${\bf x}_{t}]$ with $U\subseteq[{\bf x}_{t}]$.
  • (iii)
    Suppose that ${\bf x}_{s^{-}}]$ is the left Bongartz completion and ${\bf x}_{s^{+}}]$ is the right Bongartz completion of $U$ with respect to vertex $t_{0}$. Let ${\bf x}_{t}]$ be a cluster of $\mathcal{A}$. Then $U\subseteq[{\bf x}_{t}]$ if and only if
    $\operatorname{\mathsf{dom}}^{t_{0}}[s^{-}]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[t]\subseteq\operatorname{\mathsf{dom}}^{t_{0}}[s^{+}].$
3. $\tau$-tilting theory
3.1. Decorated modules and $\tau$-tilting pairs
We fix a finite dimensional basic algebra $A$ over an algebraically closed field $\mathbf{k}$. Denote by $\operatorname{\mathsf{mod}}\nolimits A$ the category of finitely generated left $A$-modules, and by $\tau$ the Auslander-Reiten translation in $\operatorname{\mathsf{mod}}\nolimits A$. The isomorphism classes of indecomposable projective modules and simple modules in $\operatorname{\mathsf{mod}}\nolimits A$ are denoted by $P_{1},\ldots,P_{n}$.
Given two modules $M,N\in\operatorname{\mathsf{mod}}\nolimits A$, we denote by
  • $\hom_{A}(M,N):=\dim_{\mathsf{k}}\operatorname{\mathsf{Hom}}_{A}(M,N)$.
  • $|M|$ the number of non-isomorphic indecomposable direct summands of $M$.
  • $\operatorname{\mathsf{add}}\nolimits M$ the additive closure of $M$ in $\operatorname{\mathsf{mod}}\nolimits A$.
  • $\operatorname{\mathsf{Fac}}M$ the factor modules of the modules in $\operatorname{\mathsf{add}}\nolimits M$.
  • $\operatorname{\mathsf{Sub}}M$ the submodules of the modules in $\operatorname{\mathsf{add}}\nolimits M$.
  • $\prescript{\bot}{}{M}\coloneqq\{X\in\operatorname{\mathsf{mod}}\nolimits A\mid\operatorname{\mathsf{Hom}}_{A}(X,M)=0\}$.
  • $M^{\bot}\coloneqq\{Y\in\operatorname{\mathsf{mod}}\nolimits A\mid\operatorname{\mathsf{Hom}}_{A}(M,Y)=0\}$.
A pair $\mathcal{M}=(M,P)$ of modules in $\operatorname{\mathsf{mod}}\nolimits A$ is called a decorated module of $A$, if $P$ is a projective $A$-module. The modules $M$ and $P$ are respectively called the positive part and negative part of $\mathcal{M}$. A decorated $A$-module $\mathcal{M}=(M,P)$ is called negative, if $M=0$.
Remark 3.1.
Note that the negative parts of decorated modules in our definition are projective $A$-modules, whereas the negative parts of decorated modules using in [DWZ10] are semisimple $A$-modules. Clearly, these two types of decorated $A$-modules are in bijection with each other.
Let $\mathcal{M}=(M,P)$ be a decorated $A$-module. Let $M=\oplus_{i\in I}M_{i}^{a_{i}}$ and $P=\oplus_{j\in J}P_{j}^{b_{j}}$ be the indecomposable direct sum decompositions of $M$ and $P$. Then we write $\mathcal{M}=\oplus_{k\in I\sqcup J}\mathcal{M}_{k}$, where
$\mathcal{M}_{k}=\begin{cases}(M_{k},0),&\text{if }k\in I,\\ (0,P_{k}),&\text{if }k\in J.\end{cases}$
A decorated $A$-module $\mathcal{M}=(M,P)$ is called basic, if both $M$ and $P$ are basic $A$-module. The direct sum of two decorated $A$-modules $\mathcal{M}=(M,P)$ and $\mathcal{N}=(N,Q)$ are defined as follows:
$\mathcal{M}\oplus\mathcal{N}:=(M\oplus N,P\oplus Q).$
A module $M\in\operatorname{\mathsf{mod}}\nolimits A$ is $\tau$-rigid, if $\operatorname{\mathsf{Hom}}_{A}(M,\tau M)=0$.
Definition 3.2 ($\tau$-rigid pair and $\tau$-tilting pair).
Let $\mathcal{M}=(M,P)$ be a decorated $A$-module.
  • (i)
    $\mathcal{M}=(M,P)$ is called $\tau$-rigid if $M$ is $\tau$-rigid and $\operatorname{\mathsf{Hom}}_{A}(P,M)=0$.
  • (ii)
    $\mathcal{M}=(M,P)$ is called $\tau$-tilting (resp. almost $\tau$-tilting) if $\mathcal{M}=(M,P)$ is $\tau$-rigid and
    $|M|+|P|=|A|\;\;\;(\text{resp.}\;\;|M|+|P|=|A|-1).$
We always consider modules, decorated modules up to isomorphism. In a basic $\tau$-tilting pair $(M,P)$, it is known from [air_2014, Proposition 2.3] that $P$ is uniquely determined by $M$.
Theorem 3.3 ([air_2014]*Theorem 2.12 and Theorem 2.18).
Let $\mathcal{U}=(U,Q)$ be a basic $\tau$-rigid pair. Then the following statments hold.
  • (i)
    We have $\operatorname{\mathsf{Fac}}U\subseteq\prescript{\bot}{}{(\tau U)}\cap Q^{\bot}$. The equality holds if and only if $\mathcal{U}=(U,Q)$ is $\tau$-tilting.
  • (ii)
    Suppose that $\mathcal{U}=(U,Q)$ is almost $\tau$-tilting. Then $\operatorname{\mathsf{Fac}}U\subsetneq\prescript{\bot}{}{(\tau U)}\cap Q^{\bot}$ and there exist exactly two basic $\tau$-tilting pairs $\mathcal{M}=(M,P)$ and $\mathcal{M}^{\prime}=(M^{\prime},P^{\prime})$ containing $\mathcal{U}=(U,Q)$ as a direct summand. Moreover,
    $\{\operatorname{\mathsf{Fac}}M,\;\operatorname{\mathsf{Fac}}M^{\prime}\}=\{\operatorname{\mathsf{Fac}}U,\;\prescript{\bot}{}{(\tau U)}\cap Q^{\bot}\}.$
    In particular, either $\operatorname{\mathsf{Fac}}M\subsetneq\operatorname{\mathsf{Fac}}M^{\prime}$ or $\operatorname{\mathsf{Fac}}M^{\prime}\subsetneq\operatorname{\mathsf{Fac}}M$ holds.
Definition 3.4 (Left and right mutation).
Keep the notations in Theorem 3.3. The operation $(M,P)\mapsto(M^{\prime},P^{\prime})$ is called a mutation of $(M,P)$. If $\operatorname{\mathsf{Fac}}M\subsetneq\operatorname{\mathsf{Fac}}M^{\prime}$ holds, we call $(M^{\prime},P^{\prime})$ a right mutation of $(M,P)$. If $\operatorname{\mathsf{Fac}}M^{\prime}\subsetneq\operatorname{\mathsf{Fac}}M$ holds, we call $(M^{\prime},P^{\prime})$ a left mutation of $(M,P)$.
Proposition 3.5 ([air_2014, Theorem 2.35]).
Let $\mathcal{M}=(M,P)$ and $\mathcal{N}=(N,Q)$ be two basic $\tau$-tilting pairs with $\operatorname{\mathsf{Fac}}N\subsetneqq\operatorname{\mathsf{Fac}}M$. Then there exists a left mutation $\mathcal{M}^{\prime}=(M^{\prime},P^{\prime})$ of $\mathcal{M}=(M,P)$ such that $\operatorname{\mathsf{Fac}}N\subseteq\operatorname{\mathsf{Fac}}M^{\prime}\subsetneq\operatorname{\mathsf{Fac}}M$.
3.2. $g$-vectors and $F$-polynomials of decorated modules
Let $M$ be a module in $\operatorname{\mathsf{mod}}\nolimits A$ and let
$\bigoplus_{i=1}^{n}P_{i}^{b_{i}}\rightarrow\bigoplus_{i=1}^{n}P_{i}^{a_{i}}\rightarrow M\rightarrow 0$
be the minimal projective presentation of $M$ in $\operatorname{\mathsf{mod}}\nolimits A$. The vector
$\delta_{M}\coloneqq(a_{1}-b_{1},\ldots,a_{n}-b_{n})^{T}\in\mathbb{Z}^{n}$
is called the $\delta$-vector of $M$ and the vector ${\bf g}_{M}\coloneqq-\delta_{M}$ is called the $g$-vector of $M$.
For a decorated $A$-module $\mathcal{M}=(M,P)$, we define its $\delta$-vector and $g$-vector as follows:
$\delta_{\mathcal{M}}\coloneqq\delta_{M}-\delta_{P},\;\;\;{\bf g}_{\mathcal{M}}\coloneqq-\delta_{\mathcal{M}}={\bf g}_{M}-{\bf g}_{P}.$
With this definition, we can see that the $g$-vector ${\bf g}_{(0,P_{k})}$ of $(0,P_{k})$ is the $k$-th column of $I_{n}$. This also corresponds to the $g$-vector of the $k$-th initial cluster variable $x_{k}$ in cluster algebras.
Remark 3.6.
Note that the $\delta$-vectors defined here coincide with the $g$-vectors used in [air_2014]. For the considerations on the cluster algebras side, the $g$-vectors defined here are the negative of the $\delta$-vectors.
Definition 3.7 ($F$-polynomial and dual $F$-polynomial).
(i) The $F$-polynomial $F_{M}$ of a module $M\in\operatorname{\mathsf{mod}}\nolimits A$ is defined to be
$F_{M}=\sum_{{\bf v}\in\mathbb{N}^{n}}\chi(\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(M)){\bf y}^{\bf v}\in\mathbb{Z}[y_{1},\ldots,y_{n}],$
where $\operatorname{\mathsf{Gr}}\nolimits_{\bf v}(M)$ is the quotient module Grassmannian of $M$ with dimension vector $\bf v$} and $\chi$ is the Euler–Poincaré characteristic.
(ii) The dual $F$-polynomial $\check{F}_{M}$ of a module $M\in\operatorname{\mathsf{mod}}\nolimits A$ is defined to be
$\check{F}_{M}=\sum_{{\bf v}\in\mathbb{N}^{n}}\chi(\check{\operatorname{\mathsf{Gr}}}\nolimits_{\bf v}(M)){\bf y}^{\bf v}\in\mathbb{Z}[y_{1},\ldots,y_{n}],$
where $\check{\operatorname{\mathsf{Gr}}}\nolimits_{\bf v}(M)$ is the submodule Grassmannian of $M$ with dimension vector $\bf v$} and $\chi$ is the Euler–Poincaré characteristic.
(iii) The $F$-polynomial of a decorated $A$-module $\mathcal{M}=(M,P)$ is defined by $F_{\mathcal{M}}:=F_{M}$.
Since the zero module is a quotient module of $M\in\operatorname{\mathsf{mod}}\nolimits A$, the polynomial $F_{M}$ has constant term $1$. Similarly, the dual $F$-polynomial $\check{F}_{M}$ also has the constant term $1$. Thus we have
$F_{M}[{\bf r}],\;\check{F}_{M}[{\bf r}]\in\mathbb{Z}_{\geq 0},\;\;\;\forall{\bf r}\in\mathbb{Z}^{n},$
where $F_{M}[-],\check{F}_{M}[-]:\mathbb{R}^{n}\to\mathbb{R}$ are the tropical polynomials defined in Definition 2.5.
For a negative decorated $A$-module $\mathcal{M}=(0,P)$, we clearly have $F_{\mathcal{M}}=1$.
Proposition 3.8 ([DWZ10, Proposition 3.2]).
Let $M$ and $N$ be two modules in $\operatorname{\mathsf{mod}}\nolimits A$. Then
$F_{M\oplus N}=F_{M}\cdot F_{N}.$
Proposition 3.9 ([fei_2019a, Theorem 1.4]).
For each module $M\in\operatorname{\mathsf{mod}}\nolimits A$, we have $\mathsf{P}(F_{M})=\mathsf{P}(M)$, where $\mathsf{P}(F_{M})$ is the Newton polytope of the $F$-polynomial $F_{M}$ and $\mathsf{P}(M)$ is the Newton polytope of module $M$.
Corollary 3.10.
Let $M$ and $N$ be two modules in $\operatorname{\mathsf{mod}}\nolimits A$. Then $\mathsf{P}(M\oplus N)=\mathsf{P}(M)+\mathsf{P}(N)$.
Proof.
By Propositon 3.8 and Propositon 2.4, we know that
$\mathsf{P}(F_{M\oplus N})=\mathsf{P}(F_{M}F_{N})=\mathsf{P}(F_{M})+\mathsf{P}(F_{N}).$
Then the desired result follows from Proposition 3.9. ∎
Example 3.11.
Let $A$ be the path algebra of the quiver $1\rightarrow 2$. We have the exact sequence
$0\rightarrow P_{2}\rightarrow P_{1}\rightarrow S_{1}\rightarrow 0$
in $\operatorname{\mathsf{mod}}\nolimits A=\operatorname{\mathsf{add}}\nolimits(P_{1}\oplus P_{2}\oplus S_{1})$. The Hasse quiver of the poset of functorially torsion classes in $\operatorname{\mathsf{mod}}\nolimits A$ is as follows:
${\mathcal{T}_{0}=\{0\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{1}=\operatorname{\mathsf{add}}\nolimits(S_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{4}=\operatorname{\mathsf{add}}\nolimits(P_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{2}=\operatorname{\mathsf{add}}\nolimits(S_{1}\oplus P_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{3}=\operatorname{\mathsf{mod}}\nolimits A}$
The corresponding basic $\tau$-tilting pairs are given as follows:
${(0,P_{1}\oplus P_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{(0,P_{1})}}$$\scriptstyle{\mu_{(0,P_{2})}}$${(S_{1},P_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{(0,P_{2})}}$${(P_{2},P_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{(0,P_{1})}}$${(S_{1}\oplus P_{1},0)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{(S_{1},0)}}$${\mathcal{(}P_{1}\oplus P_{2},0)}$
It is easy to check that
$F_{S_{1}}=1+y_{1},\;\;F_{P_{1}}=1+y_{1}+y_{1}y_{2},\;\;F_{P_{2}}=1+y_{2},$
${\bf g}_{S_{1}}=\begin{bmatrix}-1\\ 1\end{bmatrix},\;\;{\bf g}_{P_{1}}=\begin{bmatrix}-1\\ 0\end{bmatrix},\;\;{\bf g}_{P_{2}}=\begin{bmatrix}0\\ -1\end{bmatrix}$.
From the viewpoint of categorification of cluster algebras, the indecomposable $\tau$-rigid modules $S_{1},P_{1},P_{2}$ correspond to the non-initial cluster variables
$x_{3}={x_{1}^{-1}x_{2}\cdot(1+\widehat{y}_{1}),}\;\;\;x_{4}={x_{1}^{-1}\cdot(1+\widehat{y}_{1}+\widehat{y}_{1}\widehat{y}_{2}),}\;\;\;x_{5}={x_{2}^{-1}\cdot(1+\widehat{y}_{2})}$
in Example 2.16.
3.3. Bongartz completion in $\tau$-tilting theory
A torsion pair $(\mathcal{T},\mathcal{F})$ in $\operatorname{\mathsf{mod}}\nolimits A$ is a pair of subcategories of $\operatorname{\mathsf{mod}}\nolimits A$ satisfying that
  • (i)
    $\operatorname{\mathsf{Hom}}_{A}(T,F)=0$ for any $T\in\mathcal{T}$ and $F\in\mathcal{F}$;
  • (ii)
    for any object $X\in\operatorname{\mathsf{mod}}\nolimits A$, there exists a short exact sequence
    $0\rightarrow X_{t}\rightarrow X\rightarrow X_{f}\rightarrow 0$
    with $X_{t}\in\mathcal{T}$ and $X_{f}\in\mathcal{F}$. Thanks to the condition (i), such a sequence is unique up to isomorphisms. This short exact sequence is called the canonical sequence of $X$ with respect to $(\mathcal{T},\mathcal{F})$.
Notice that in a torsion pair $(\mathcal{T},\mathcal{F})$, we always have $\mathcal{F}=\mathcal{T}^{\bot}$ and $\mathcal{T}=\prescript{\bot}{}{\mathcal{F}}$. The subcategory $\mathcal{T}$ (resp., $\mathcal{F}$) in a torsion pair $(\mathcal{T},\mathcal{F})$ is called a torsion class (resp., torsion-free class) in $\operatorname{\mathsf{mod}}\nolimits A$.
It is known that a subcategory of $\operatorname{\mathsf{mod}}\nolimits A$ is a torsion class if and only if it is closed under extensions and quotients. A torsion class $\mathcal{T}$ is said to be functorially finite, if there exists a module $M\in\operatorname{\mathsf{mod}}\nolimits A$ such that $\mathcal{T}=\operatorname{\mathsf{Fac}}M$.
A module $M$ in a subcategory $\mathcal{C}$ of $\operatorname{\mathsf{mod}}\nolimits A$ is said to be Ext-projective in $\mathcal{C}$, if $\operatorname{\mathsf{Ext}}_{A}^{1}(M,\mathcal{C})=0$. We denote by $\mathcal{P}(\mathcal{C})$ the direct sum of one copy of each of the indecomposable Ext-projective objects in $\mathcal{C}$ up to isomorphism.
Theorem 3.12 ([air_2014, Proposition 1.2, Theorem 2.7]).
The following statements hold.
  • (i)
    There is a map $\Psi$ from $\tau$-rigid pairs to functorially finite torsion classes in $\operatorname{\mathsf{mod}}\nolimits A$ given by $(M,P)\mapsto\operatorname{\mathsf{Fac}}M$.
  • (ii)
    The above map $\Psi$ is a bijection if we restrict it to basic $\tau$-tilting pairs.
  • (iii)
    Let $(M,P)$ be a basic $\tau$-tilting pair. Then $M=\mathcal{P}(\operatorname{\mathsf{Fac}}M)$.
Proposition 3.13 ([air_2014]*Proposition 2.9, Theorem 2.10).
Let $(U,Q)$ be a basic $\tau$-rigid pair in $\operatorname{\mathsf{mod}}\nolimits A$. Then
  • (i)
    $\operatorname{\mathsf{Fac}}U$ and $\prescript{\bot}{}{(}{\tau U})\cap Q^{\bot}$ are functorially finite torsion classes in $\operatorname{\mathsf{mod}}\nolimits A$.
  • (ii)
    $(U,Q)$ is a direct summand of a $\tau$-tilting pair $(M,P)$ if and only if
    $\operatorname{\mathsf{Fac}}U\subseteq\operatorname{\mathsf{Fac}}M\subseteq\prescript{\bot}{}{(}{\tau U})\cap Q^{\bot}.$
Definition 3.14 (Left and right Bongartz completion).
Let $(U,Q)$ be a basic $\tau$-rigid pair in $\operatorname{\mathsf{mod}}\nolimits A$ and $\Psi$ the bijection in Theorem 3.12 (ii).
  • (i)
    The left Bongartz completion (or Bongartz co-completion) of $(U,Q)$ is defined to be the basic $\tau$-tilting pair $(M^{-},P^{-})$ such that $\operatorname{\mathsf{Fac}}M^{-}=\operatorname{\mathsf{Fac}}U$.
  • (ii)
    The right Bongartz completion (or Bongartz completion) of $(U,Q)$ is defined to be the basic $\tau$-tilting pair $(M^{+},P^{+})$ such that $\operatorname{\mathsf{Fac}}M^{+}=\prescript{\bot}{}{(}{\tau U})\cap Q^{\bot}$.
Remark 3.15.
It is each to see that the left Bongartz completion $(M^{-},P^{-})$ of $(U,Q)$ can be characterized by the following two conditions:
  • (a)
    $(U,Q)$ is a direct summand of $(M^{-},P^{-})$;
  • (b)
    $\operatorname{\mathsf{Fac}}M^{-}\subseteq\operatorname{\mathsf{Fac}}M$ for any basic $\tau$-tilting pair $(M,P)$ such that $(U,Q)$ is a direct summand of $(M,P)$.
The right Bongart completion $(M^{+},P^{+})$ of $(U,Q)$ can be characterized in a similar way.
4. $F$-invariant and Newton polytopes of $F$-polynomials
4.1. $F$-invariant of cluster monomials
$F$-invariant in cluster algebras is introduced by the author in [Cao-2023], which generalizes the $E$-invariant [DWZ10] in additive categorification of cluster algebras and the $\mathfrak{d}$-invariant [kkko-2018, kkop-2020] in monoidal categorification of cluster algebras.
Recall that for a nonzero polynomial $F({\bf y})=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}{\bf y}^{\bf v}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$, its tropical polynomial $F[-]:\mathbb{R}^{n}\to\mathbb{R}$ is defined by
$F[{\bf r}]:=\max\{\langle{{\bf v},{\bf r}}\rangle\mid c_{\bf v}\neq 0\}.$
Definition 4.1 ($F$-invariant and partial $F$-invariant).
Let $\mathcal{A}$ be a cluster algebra and $D=diag(d_{1},\ldots,d_{n})$ a fixed skew-symmetrizer for the exchange matrices of $\mathcal{A}$. Let $u$ and $v$ be two cluster monomials of $\mathcal{A}$, and let
$u={\bf x}_{t}^{{\bf g}_{u}^{t}}F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t})\;\;\;\;\text{and}\;\;\;v={\bf x}_{t}^{{\bf g}_{v}^{t}}F_{v}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t})$
be the canonical expressions of $u$ and $v$ with repect to vertex $t\in\mathbb{T}_{n}$.
  • (i)
    The integer
    $(u\mid\mid v)_{F}=F_{u}^{t}[D{\bf g}_{v}^{t}]+F_{v}^{t}[D{\bf g}_{u}^{t}]$
    is called the $F$-invariant of $(u,v)$, which is independent of the choice of $t\in\mathbb{T}_{n}$ by Theorem 4.2 below.
  • (ii)
    The integer $F_{u}^{t}[D{\bf g}_{v}^{t}]$ is called the partial $F$-invariant of $(u,v)$ at vertex $t\in\mathbb{T}_{n}$.
Since the definition of $F$-invariant depends on the choice of a skew-symmetrizer $D$, we always fix a skew-symmetrizer for the exchange matrices of $\mathcal{A}$ in this paper.
Theorem 4.2 ([Cao-2023, Theorem 4.10, Proposition 4.30]).
Let $u$ and $v$ be two cluster monomials of $\mathcal{A}$ and $D=diag(d_{1},\ldots,d_{n})$ a fixed skew-symmetrizer for the exchange matrices of $\mathcal{A}$. Then for any two vertices $t,t^{\prime}\in\mathbb{T}_{n}$, we have
$F_{u}^{t}[D{\bf g}_{v}^{t}]+F_{v}^{t}[D{\bf g}_{u}^{t}]=F_{u}^{t^{\prime}}[D{\bf g}_{v}^{t^{\prime}}]+F_{v}^{t^{\prime}}[D{\bf g}_{u}^{t^{\prime}}].$
In particular, the $F$-invariant $(u\mid\mid v)_{F}=F_{u}^{t}[D{\bf g}_{v}^{t}]+F_{v}^{t}[D{\bf g}_{u}^{t}]$ only depends on $u$ and $u^{\prime}$, not on the choice of vertex $t\in\mathbb{T}_{n}$.
Since the $F$-polynomials of cluster monomials have constant term $1$, we have
$(u\mid\mid v)_{F}=F_{u}^{t}[D{\bf g}_{v}^{t}]+F_{v}^{t}[D{\bf g}_{u}^{t}]\geq 0.$
Theorem 4.3 ([Cao-2023, Theorem 4.19]).
Let $u$ and $v$ be two cluster monomials of $\mathcal{A}$. Then the product $u\cdot v$ remains a cluster monomial of $\mathcal{A}$ if and only if $(u\mid\mid v)_{F}=0$.
Example 4.4.
Let us continue the example of cluster algebra $\mathcal{A}$ of type $A_{2}$ in Example 2.16. We know that the canonical expressions of $x_{3},x_{4},x_{5}$ with respect to $t_{0}=({\bf x},B)$ are given as follows:
$x_{3}={x_{1}^{-1}x_{2}\cdot(1+\widehat{y}_{1}),}\;\;\;x_{4}={x_{1}^{-1}\cdot(1+\widehat{y}_{1}+\widehat{y}_{1}\widehat{y}_{2}),}\;\;\;x_{5}={x_{2}^{-1}\cdot(1+\widehat{y}_{2})}.$
We take $D=I_{2}$ to be the fixed skew-symmetrizer for the exchange matrices of $\mathcal{A}$. We have
$(x_{3}\mid\mid x_{4})_{F}$ $=F_{x_{3}}^{t_{0}}[D{\bf g}_{x_{4}}^{t_{0}}]+F_{x_{4}}^{t_{0}}[D{\bf g}_{x_{3}}^{t_{0}}]$
$=(1+y_{1})\begin{bmatrix}-1\\ 0\end{bmatrix}+(1+y_{1}+y_{1}y_{2})\begin{bmatrix}-1\\ 1\end{bmatrix}$
$=\max\{0,\;-1\}+\max\{0,\;-1,\;0\}$
$=0,$
$(x_{3}\mid\mid x_{5})_{F}$ $=(1+y_{1})\begin{bmatrix}0\\ -1\end{bmatrix}+(1+y_{2})\begin{bmatrix}-1\\ 1\end{bmatrix}$
$=\max\{0,\;0\}+\max\{0,\;1\}$
$=1.$
Then by Theorem 4.3, we know that $x_{3}x_{4}$ is a cluster monomial, while $x_{3}x_{5}$ is not.
Proposition 4.5 ([Cao-2023, Propositions 4.9, 4.16]).
Let $u={\bf x}_{t}^{\bf h}=\prod_{i=1}^{n}x_{i;t}^{h_{i}}$ and $v$ be two cluster monomials of $\mathcal{A}$. Then we have $(u\mid\mid v)_{F}=\sum_{i=1}^{n}h_{i}(x_{i;t}\mid\mid v)_{F}$. Moreover, we have
$F_{v}^{w}[D{\bf g}_{u}^{w}]=\sum_{i=1}^{n}h_{i}F_{v}^{w}[D{\bf g}_{x_{i;t}}^{w}]\;\;\;\text{and}\;\;\;F_{u}^{w}[D{\bf g}_{v}^{w}]=\sum_{i=1}^{n}h_{i}F_{x_{i;t}}^{w}[D{\bf g}_{v}^{w}]$
for any vertex $w\in\mathbb{T}_{n}$.
4.2. Newton polytopes of $F$-polynomials
Lemma 4.6.
Let $\mathcal{A}$ be a cluster algebra with initial seed $({\bf x}_{t_{0}},B_{t_{0}})$. If two cluster variables $x$ and $z$ are contained in the same cluster, then $F_{z}^{t_{0}}[D{\bf g}_{x}^{t_{0}}]=0=F_{x}^{t_{0}}[D{\bf g}_{z}^{t_{0}}]$.
Proof.
Since the two cluster variables $x$ and $z$ are contained in the same cluster and by Theorem 4.3, we have $F_{x}^{t_{0}}[D{\bf g}_{z}^{t_{0}}]+F_{z}^{t_{0}}[D{\bf g}_{x}^{t_{0}}]=(x\mid\mid z)_{F}=0$. Since $F_{x}^{t_{0}}[D{\bf g}_{z}^{t_{0}}]\geq 0$ and $F_{z}^{t_{0}}[D{\bf g}_{x}^{t_{0}}]\geq 0$, we get $F_{z}^{t_{0}}[D{\bf g}_{x}^{t_{0}}]=0=F_{x}^{t_{0}}[D{\bf g}_{z}^{t_{0}}]$. ∎
Proposition 4.7.
Let $\mathcal{A}$ be a cluster algebra with initial seed $({\bf x}_{t_{0}},B_{t_{0}})$ and $D=diag(d_{1},\ldots,d_{n})$ a fixed skew-symmetrizer for the exchange matrices of $\mathcal{A}$. Let $({\bf x}_{t^{\prime}},B_{t^{\prime}})=\mu_{k}({\bf x}_{t},B_{t})$ be a mutation in $\mathcal{A}$. Then $(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=d_{k}$. If moreover, the $k$-th column of $C_{t}^{t_{0}}$ is non-positive, then we have
$F_{x_{k;t^{\prime}}}^{t_{0}}[{\bf g}_{x_{k;t}}]=0\quad\text{and}\quad F_{x_{k;t}}^{t_{0}}[{\bf g}_{x_{k;t^{\prime}}}]=d_{k}.$
Proof.
Let us use the vertex $t$ to calculate the $F$-invariant $(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}$. Since $F_{x_{k;t}}^{t}=1,F_{x_{k;t^{\prime}}}^{t}=1+y_{k}$ and ${\bf g}_{x_{k;t}}^{t}={\bf e}_{k}$, We have
$(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=F_{x_{k;t}}^{t}[D{\bf g}_{k;t^{\prime}}^{t}]+F_{x_{k;t^{\prime}}}^{t}[D{\bf g}_{x_{k;t}}^{t}]=0+(1+y_{k})[D{\bf e}_{k}]=d_{k}.$
On the other hand, if we use the vertex $t_{0}$ to calculate $(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}$, we have
$d_{k}=(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=F_{x_{k;t}}^{t_{0}}[D{\bf g}_{k;t^{\prime}}^{t}]+F_{x_{k;t^{\prime}}}^{t_{0}}[D{\bf g}_{x_{k;t}}^{t}].$
Now let us prove that $F_{x_{k;t^{\prime}}}^{t_{0}}[D{\bf g}_{k;t}^{t}]=0$ under the assumption that the $k$-th column ${\bf c}_{k;t}$ of $C_{t}^{t_{0}}$ is non-positive. In this case, we have ${\bf c}_{k;t}\in\mathbb{Z}_{\geq 0}^{n}$. By the mutation rule [fomin_zelevinsky_2007, Proposition 5.1] of $F$-polynomials, we have
$F_{x_{k;t}}^{t_{0}}\cdot F_{x_{k;t^{\prime}}}^{t_{0}}=\prod_{b_{ik}^{t}>0}(F_{x_{i;t}}^{t_{0}})^{b_{ik}^{t}}+{\bf y}^{-{\bf c}_{k;t}}\cdot\prod_{b_{ik}^{t}<0}(F_{x_{i;t}}^{t_{0}})^{-b_{ik}^{t}}.$
By Corollary 2.9, we have the tropical version of the above equality:
(4.1) $F_{x_{k;t}}^{t_{0}}[{\bf r}]+F_{x_{k;t^{\prime}}}^{t_{0}}[{\bf r}]=\max\{\sum_{b_{ik}^{t}>0}b_{ik}^{t}F_{x_{i;t}}^{t_{0}}[{\bf r}],\;\langle-{\bf c}_{k;t},{\bf r}\rangle+\sum_{b_{ik}^{t}<0}(-b_{ik}^{t})F_{x_{i;t}}^{t_{0}}[{\bf r}]\}$
for any ${\bf r}\in\mathbb{R}^{n}$. By Lemma 4.6, we know that $F_{x_{i;t}}^{t_{0}}[D{\bf g}_{k;t}^{t_{0}}]=0$ for any $i\in[1,n]$. Then by taking ${\bf r}=D{\bf g}_{x_{k;t}}^{t_{0}}$ in (4.1), we get
$0+F_{x_{k;t^{\prime}}}^{t_{0}}[D{\bf g}_{x_{k;t}}^{t_{0}}]=\max\{0,\;\langle-{\bf c}_{k;t},D{\bf g}_{x_{k;t}}^{t_{0}}\rangle+0\}=\max\{0,\;\langle-{\bf c}_{k;t},D{\bf g}_{x_{k;t}}^{t_{0}}\rangle\}.$
By Theorem 2.18 (i), we have $(G_{t}^{t_{0}})^{T}DC_{t}^{t_{0}}D^{-1}=I_{n}$. Thus $\langle-{\bf c}_{k;t},D{\bf g}_{x_{k;t}}^{t_{0}}\rangle\}=-d_{k}$. Hence, we have
$F_{x_{k;t^{\prime}}}^{t_{0}}[D{\bf g}_{x_{k;t}}^{t_{0}}]=\max\{0,\;\langle-{\bf c}_{k;t},D{\bf g}_{x_{k;t}}^{t_{0}}\rangle\}=\max\{0,\;-d_{k}\}=0.$
Thus $F_{x_{k;t}}^{t_{0}}[{\bf g}_{x_{k;t^{\prime}}}^{t_{0}}]=(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}-F_{x_{k;t^{\prime}}}^{t_{0}}[D{\bf g}_{x_{k;t}}^{t_{0}}]=d_{k}-0=d_{k}$. ∎
Recall that for a non-zero polynomial $F\in\mathbb{Z}[y_{1},\ldots,y_{n}]$, $\mathsf{P}(F)$ denotes the Newton polytope of $F$.
Lemma 4.8.
Let $u=\prod_{i=1}^{n}x_{i;t}^{a_{i}}$ and $v=\prod_{i=1}^{n}x_{i;t}^{b_{i}}$ be two cluster monomials of $\mathcal{A}$ in some cluster ${\bf x}_{t}$. If $\mathsf{P}(F_{u}^{t_{0}})=\mathsf{P}(F_{v}^{t_{0}})$, then $a_{k}=b_{k}$ whenever the $k$-th column of the $C$-matrix $C_{t}^{t_{0}}$ is non-positive.
Proof.
Suppose that $k$-th column of the $C$-matrix $C_{t}^{t_{0}}$ is non-positive. By Proposition 4.7, we have
$F_{x_{k;t}}^{t_{0}}[D{\bf g}_{x_{k;t^{\prime}}}^{t_{0}}]=d_{k}>0.$
For $i\neq k$, we know that $\{x_{i;t},x_{k;t^{\prime}}\}\subseteq[{\bf x}_{t^{\prime}}]$. Then by Lemma 4.6, we have
(4.2) $F_{x_{i;t}}^{t_{0}}[D{\bf g}_{x_{k;t^{\prime}}}^{t_{0}}]=0,\quad\forall i\neq k.$
Since $u=\prod_{i=1}^{n}x_{i;t}^{a_{i}}$ and $v=\prod_{i=1}^{n}x_{i;t}^{b_{i}}$ and by Proposition 4.5, we have
[…]
Since the two $F$-polynomials $F_{u}^{t_{0}}$ and $F_{v}^{t_{0}}$ have the same Newton polytope, we know that $F_{u}^{t_{0}}[{\bf r}]=F_{v}^{t_{0}}[{\bf r}]$ for any vector ${\bf r}\in\mathbb{Z}^{n}$. In particular, we have
$a_{k}d_{k}=F_{u}^{t_{0}}[D{\bf g}_{k;t^{\prime}}^{t_{0}}]=F_{v}^{t_{0}}[D{\bf g}_{k;t^{\prime}}^{t_{0}}]=b_{k}d_{k}.$
Since $d_{k}>0$, we obtain $a_{k}=b_{k}$. ∎
Lemma 4.9.
Let $u=\prod_{i=1}^{n}x_{i;t}^{a_{i}}$ and $v=\prod_{i=1}^{n}x_{i;t}^{b_{i}}$ be two cluster monomials of $\mathcal{A}$ in some cluster ${\bf x}_{t}$. If $\mathsf{P}(F_{u}^{t_{0}})=\mathsf{P}(F_{v}^{t_{0}})$ and $a_{k}=b_{k}$ for some $k\in[1,n]$, then $\mathsf{P}(F_{u^{\prime}}^{t_{0}})=\mathsf{P}(F_{v^{\prime}}^{t_{0}})$, where $u^{\prime}=u/x_{k;t}^{a_{k}}=\prod_{i\neq k}^{n}x_{i;t}^{a_{i}}$ and $v^{\prime}=v/x_{k;t}^{b_{k}}=\prod_{i\neq k}^{n}x_{i;t}^{b_{i}}$.
Proof.
We know that $u=u^{\prime}\cdot x_{k;t}^{a_{k}}$ and $v=v^{\prime}\cdot x_{k;t}^{b_{k}}=v^{\prime}\cdot x_{k;t}^{a_{k}}$. Denote by $H=(F_{x_{k;t}}^{t_{0}})^{a_{k}}$. Then we have $F_{u}^{t_{0}}=F_{u^{\prime}}^{t_{0}}\cdot H$ and $F_{v}^{t_{0}}=F_{v^{\prime}}^{t_{0}}\cdot H$. By Proposition 2.4, we have $\mathsf{P}(F_{u}^{t_{0}})=\mathsf{P}(F_{u^{\prime}}^{t_{0}})+\mathsf{P}(H)$ and $\mathsf{P}(F_{v}^{t_{0}})=\mathsf{P}(F_{v^{\prime}}^{t_{0}})+\mathsf{P}(H)$. Since $\mathsf{P}(F_{u}^{t_{0}})=\mathsf{P}(F_{v}^{t_{0}})$ and by Corollary 2.2, we get $\mathsf{P}(F_{u^{\prime}}^{t_{0}})=\mathsf{P}(F_{v^{\prime}}^{t_{0}})$. ∎
Lemma 4.10 ([Cao-2023, Corollary 4.21]).
Let $u$ and $v$ be two cluster monomials of $\mathcal{A}$. If $\mathsf{P}(F_{u}^{t_{0}})=\mathsf{P}(F_{v}^{t_{0}})$, then the product $uv$ remains a cluster monomials of $\mathcal{A}$.
Proof.
Since $u$ and $v$ are cluster monomials and by Theorem 4.3, we have
$(u\mid\mid u)_{F}=2F_{u}^{t_{0}}[D{\bf g}_{u}^{t_{0}}]=0\;\;\;\text{and}\;\;\;(v\mid\mid v)_{F}=2F_{v}^{t_{0}}[D{\bf g}_{v}^{t_{0}}]=0.$
We get $F_{u}^{t_{0}}[D{\bf g}_{u}^{t_{0}}]=0=F_{v}^{t_{0}}[D{\bf g}_{v}^{t_{0}}]$. Since $F_{u}^{t_{0}}$ and $F_{v}^{t_{0}}$ have the same Newton polytope, we have $F_{u}^{t_{0}}[{\bf r}]=F_{v}^{t_{0}}[{\bf r}]$ for any ${\bf r}\in\mathbb{Z}^{n}$. Hence, we have
$(u\mid\mid v)_{F}=F_{u}^{t_{0}}[D{\bf g}_{v}^{t_{0}}]+F_{v}^{t_{0}}[D{\bf g}_{u}^{t_{0}}]=F_{v}^{t_{0}}[D{\bf g}_{v}^{t_{0}}]+F_{u}^{t_{0}}[D{\bf g}_{u}^{t_{0}}]=0.$
Then by Theorem 4.3, we know that the product uvuv is a cluster monomial. ∎
Denote by $\mathcal{X}$ the set of cluster variables of $\mathcal{A}$. Given a cluster monomial $u={\bf x}_{t}^{\bf h}=\prod x_{i;t}^{h_{i}}$ in a seed $({\bf x}_{t},B_{t})$, we define its support set $\operatorname{\mathsf{supp}}(u)$ by
$\operatorname{\mathsf{supp}}(u):=\{x_{i;t}\mid i\in[1,n],\;h_{i}\neq 0\}\subseteq\mathcal{X}.$
We remark that this set only depends on $u$, not on the choice of $t\in\mathbb{T}_{n}$ and ${\bf h}\in\mathbb{Z}^{n}$ such that $u={\bf x}_{t}^{\bf h}$.
Theorem 4.11.
Let $\mathcal{A}$ be a skew-symmetrizable cluster algebra with initial seed $({\bf x}_{t_{0}},B_{t_{0}})$. Let $u$ and $v$ be two cluster monomials in non-initial cluster variables. If the two $F$-polynomials $F_{u}^{t_{0}}$ and $F_{v}^{t_{0}}$ have the same Newton polytope, i.e., $\mathsf{P}(F_{u}^{t_{0}})=\mathsf{P}(F_{v}^{t_{0}})$, then $u=v$.
Proof.
Since $\mathsf{P}(F_{u}^{t_{0}})=\mathsf{P}(F_{v}^{t_{0}})$ and by Lemma 4.10, the product $uv$ remains a cluster monomial. Thus its support set $\operatorname{\mathsf{supp}}(uv)$ is a partial cluster of $\mathcal{A}$.
Let ${\bf x}_{s}]$ be the left Bongartz completion of $\operatorname{\mathsf{supp}}(uv)$ with respect to vertex $t_{0}$. Then $\operatorname{\mathsf{supp}}(uv)\subseteq[{\bf x}_{s}]$. We can write $u=\prod_{i=1}^{n}x_{i;s}^{a_{i}}$ and $v=\prod_{i=1}^{n}x_{i;s}^{b_{i}}$. Thus $uv=\prod_{i=1}^{n}x_{i;s}^{a_{i}+b_{i}}$. We know that $a_{i}+b_{i}>0$ if and only if $x_{i;s}\in\operatorname{\mathsf{supp}}(uv)$.
Since ${\bf x}_{s}]$ is the left Bongartz completion of $\operatorname{\mathsf{supp}}(uv)$ with respect to vertex $t_{0}$, we know that the $j$-th column of the $C$-matrix $C_{s}^{t_{0}}$ is non-negative for any $j$ with $x_{j;s}\in[{\bf x}_{s}]\setminus\operatorname{\mathsf{supp}}(uv)$. For the signs of the other columns of $C_{s}^{t_{0}}$, we have the following claim.
Claim: there exist some $k$ with $x_{k;s}\in\operatorname{\mathsf{supp}}(uv)$ such that the $k$-th column of $C_{s}^{t_{0}}$ is non-positive. Otherwise, we know that $C_{s}^{t_{0}}$ is a non-negative matrix. Then by Proposition 2.20, we have ${\bf x}_{s}]=[{\bf x}_{t_{0}}]$. Thus $\operatorname{\mathsf{supp}}(uv)\subseteq[{\bf x}_{s}]=[{\bf x}_{t_{0}}]$. This contradicts that $u$ and $v$ are cluster monomials in non-initial cluster variables. This finishes the proof of the claim.
Now we prove $u=v$ by reducing the number $|\operatorname{\mathsf{supp}}(uv)|$ of cluster variables in $\operatorname{\mathsf{supp}}(uv)$. By the claim, we can take an integer $k$ with $x_{k;s}\in\operatorname{\mathsf{supp}}(uv)$ such that $k$-th column of $C_{s}^{t_{0}}$ is non-positive. Since $\mathsf{P}(F_{u}^{t_{0}})=\mathsf{P}(F_{v}^{t_{0}})$ and by Lemma 4.8, we know that $a_{k}=b_{k}$. Then by Lemma 4.9, we have
$\mathsf{P}(F_{u^{\prime}}^{t_{0}})=\mathsf{P}(F_{v^{\prime}}^{t_{0}}),$
where $u^{\prime}=u/x_{k;s}^{a_{k}}=\prod_{i\neq k}x_{i;s}^{a_{i}}$ and $v^{\prime}=v/x_{k;s}^{b_{k}}=\prod_{i\neq k}x_{i;s}^{b_{i}}$. Notice that we have
$u=u^{\prime}\cdot x_{k;s}^{a_{k}},\;\;\;v=v^{\prime}\cdot x_{k;s}^{b_{k}}=v^{\prime}\cdot x_{k;s}^{a_{k}}\;\;\;\text{and}\;\;\;\operatorname{\mathsf{supp}}(u^{\prime}v^{\prime})\subsetneq\operatorname{\mathsf{supp}}(uv).$
Thus to show $u=v$, it suffices to show $u^{\prime}=v^{\prime}$. Now $u^{\prime}$ and $v^{\prime}$ are two cluster monomials in non-initial cluster variables satisfying $\mathsf{P}(F_{u^{\prime}}^{t_{0}})=\mathsf{P}(F_{v^{\prime}}^{t_{0}})$ and $|\operatorname{\mathsf{supp}}(u^{\prime}v^{\prime})|<|\operatorname{\mathsf{supp}}(uv)|$. We can therefore apply the reduction and complete the proof. ∎
5. $F$-invariant and Newton polytopes of $\tau$-rigid modules
5.1. $F$-invariant of decorated modules
Recall that for two modules $M,N\in\operatorname{\mathsf{mod}}\nolimits A$, we denote by $\hom_{A}(M,N)=\dim_{\mathsf{k}}\operatorname{\mathsf{Hom}}_{A}(M,N)$.
Definition 5.1 ($E$-invariant and partial $E$-invariant, [DWZ10, DK-2015, air_2014]).
Let $\mathcal{M}=(M,P)$ and $\mathcal{N}=(N,Q)$ be two decorated $A$-modules. The partial $E$-invariant $E^{\rm proj}(\mathcal{M},\mathcal{N})$ and the $E$-invariant $E^{\rm sym}(\mathcal{M},\mathcal{N})$ of the ordered pair $(\mathcal{M},\mathcal{N})$ are defined as follows:
(5.1) $E^{\rm proj}(\mathcal{M},\mathcal{N})$ $:=$ $\hom_{A}(N,\tau M)+\hom_{A}(P,N),$
$E^{\rm sym}(\mathcal{M},\mathcal{N})$ $:=$ $E^{\rm proj}(\mathcal{M},\mathcal{N})+E^{\rm proj}(\mathcal{N},\mathcal{M})$
$=$ $\hom_{A}(N,\tau M)+\hom_{A}(P,N)+\hom_{A}(M,\tau N)+\hom_{A}(Q,M).$
Definition 5.2 ($F$-invariant and partial $F$-invariant).
Let $\mathcal{M}=(M,P)$ and $\mathcal{N}=(N,Q)$ be two decorated $A$-modules. Let $F_{\mathcal{M}}=F_{M}=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}{\bf y}^{\bf v}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ be the $F$-polynomial of $\mathcal{M}$ and ${\bf g}_{\mathcal{N}}\in\mathbb{Z}^{n}$ the $g$-vector of $\mathcal{N}$. The partial $F$-invariant $F_{\mathcal{M}}[{\bf g}_{\mathcal{N}}]$ and the $F$-invariant $(\mathcal{M}\mid\mid\mathcal{N})_{F}$ of the ordered pair $(\mathcal{M},\mathcal{N})$ are defined as follows:
(5.3) $F_{\mathcal{M}}[{\bf g}_{\mathcal{N}}]$ $:=$ $\max\{\langle{\bf v},{\bf g}_{\mathcal{N}}\rangle\mid c_{\bf v}\neq 0\}=\max\{\langle{\bf v},{\bf g}_{\mathcal{N}}\rangle\mid{\bf v}\in\mathsf{P}(M)\},$
(5.4) $(\mathcal{M}\mid\mid\mathcal{N})_{F}$ $:=$ $F_{\mathcal{M}}[{\bf g}_{\mathcal{N}}]+F_{\mathcal{N}}[{\bf g}_{\mathcal{M}}].$
Lemma 5.3 ([air_2014, Prop. 2.4], [AR-1985, Thm. 1.4]).
Let $M$ and $X$ be two modules in $\operatorname{\mathsf{mod}}\nolimits A$. Then [^1] Note that $g$-vectors of modules in this paper are the negative of the $g$-vectors used in [air_2014].
$\langle{\bf g}_{M},\underline{\dim}X\rangle=\hom_{A}(X,\tau M)-\hom_{A}(M,X).$
Lemma 5.4.
Let $M\in\operatorname{\mathsf{mod}}\nolimits A$ be a $\tau$-rigid module and let
(5.7) ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{t}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{f}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0,}$
be the canonical sequence of a module $N$ with respect to the torsion pair $(\operatorname{\mathsf{Fac}}M,M^{\bot})$. Then for any quotient module $N_{0}$ of $N$, we have
$\langle{\bf g}_{M},\underline{\dim}N_{0}\rangle\leq\langle{\bf g}_{M},\underline{\dim}N_{f}\rangle=\hom(N,\tau M).$
Proof.
We know that $N_{t}\in\operatorname{\mathsf{Fac}}M$ and $N_{f}\in M^{\bot}$. Since $N_{f}\in M^{\bot}$ and by Lemma 5.3, we have
$\langle{\bf g}_{M},\underline{\dim}N_{f}\rangle=\hom_{A}(N_{f},\tau M)-\hom_{A}(M,N_{f})=\hom_{A}(N_{f},\tau M).$
Since $N_{t}\in\operatorname{\mathsf{Fac}}M\subseteq\prescript{\bot}{}{(}\tau M)$ and by applying the functor $\operatorname{\mathsf{Hom}}_{A}(-,\tau M)$ to (5.7), we get $\operatorname{\mathsf{Hom}}_{A}(N_{f},\tau M)\cong\operatorname{\mathsf{Hom}}_{A}(N,\tau M)$. Hence, we have
$\langle{\bf g}_{M},\underline{\dim}N_{f}\rangle=\hom_{A}(N_{f},\tau M)=\hom_{A}(N,\tau M).$
Suppose that $N_{0}$ is a quotient module of $N$, i.e., we have an exact sequence $N\to N_{0}\to 0$. By applying the functor $\operatorname{\mathsf{Hom}}_{A}(-,\tau M)$, we see that
$\hom(N_{0},\tau M)\leq\hom(N,\tau M).$
By Lemma 5.3, we have
$\langle{\bf g}_{M},\underline{\dim}N_{0}\rangle$ $=$ $\hom_{A}(N_{0},\tau M)-\hom_{A}(M,N_{0})$
$\leq$ $\hom_{A}(N_{0},\tau M)$
$\leq$ $\hom_{A}(N,\tau M)=\langle{\bf g}_{M},\underline{\dim}N_{f}\rangle.$
This completes the proof. ∎
Lemma 5.5.
Let $M\in\operatorname{\mathsf{mod}}\nolimits A$ be a $\tau$-rigid module and let
(5.10) ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{t}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{f}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0,}$
be the canonical sequence of a module $N$ with respect to the torsion pair $(\prescript{\bot}{}{(\tau M)},\operatorname{\mathsf{Sub}}\tau M)$. Then for any submodule $N_{0}$ of $N$, we have
$\langle-{\bf g}_{M},\underline{\dim}N_{0}\rangle\leq\langle-{\bf g}_{M},\underline{\dim}N_{t}\rangle=\hom(M,N).$
Proof.
We know that $N_{t}\in\prescript{\bot}{}{(\tau M)}$ and $N_{f}\in\operatorname{\mathsf{Sub}}\tau M$. Since $N_{t}\in\prescript{\bot}{}{(\tau M)}$ and by Lemma 5.3, we have
$\langle{-\bf g}_{M},\underline{\dim}N_{t}\rangle=-\hom_{A}(N_{t},\tau M)+\hom_{A}(M,N_{t})=\hom_{A}(M,N_{t}).$
Since $\operatorname{\mathsf{Hom}}_{A}(M,\tau M)=0$, $N_{f}\in\operatorname{\mathsf{Sub}}\tau M$ and by applying the functor $\operatorname{\mathsf{Hom}}_{A}(M,-)$ to (5.10), we get $\operatorname{\mathsf{Hom}}_{A}(M,N_{t})\cong\operatorname{\mathsf{Hom}}_{A}(M,N)$. Hence, we have
$\langle-{\bf g}_{M},\underline{\dim}N_{t}\rangle=\hom_{A}(M,N_{t})=\hom_{A}(M,N).$
Suppose that $N_{0}$ is a submodule of $N$, i.e., we have an exact sequence $0\to N_{0}\to N$. By applying the functor $\operatorname{\mathsf{Hom}}_{A}(M,-)$, we see that
$\hom(M,N_{0})\leq\hom(M,N).$
By Lemma 5.3, we have
$\langle-{\bf g}_{M},\underline{\dim}N_{0}\rangle$ $=$ $-\hom_{A}(N_{0},\tau M)+\hom_{A}(M,N_{0})$
$\leq$ $\hom_{A}(M,N_{0})$
$\leq$ $\hom_{A}(M,N)=\langle-{\bf g}_{M},\underline{\dim}N_{t}\rangle.$
This completes the proof. ∎
Proposition 5.6.
Let $M\in\operatorname{\mathsf{mod}}\nolimits A$ be a $\tau$-rigid module. Then for any module $N\in\operatorname{\mathsf{mod}}\nolimits A$, we have
$F_{N}[{\bf g}_{M}]=\hom_{A}(N,\tau M)\quad\text{and}\quad\check{F}_{N}[-{\bf g}_{M}]=\hom_{A}(M,N),$
where $\check{F}_{N}$ is the dual $F$-polynomial of $N$.
Proof.
This follows from Lemma 5.4 and Lemma 5.5. ∎
Theorem 5.7.
Let $\mathcal{M}=(M,P)$ be a $\tau$-rigid pair in $\operatorname{\mathsf{mod}}\nolimits A$. Then for any decorated $A$-module $\mathcal{N}=(N,Q)$, we have
$F_{\mathcal{N}}[{\bf g}_{\mathcal{M}}]=E^{\rm proj}(\mathcal{M},\mathcal{N})=\hom_{A}(N,\tau M)+\hom_{A}(P,N).$
Proof.
Since $F_{\mathcal{N}}=F_{N}$, it suffices to show that $F_{N}[{\bf g}_{\mathcal{M}}]=\hom_{A}(N,\tau M)+\hom_{A}(P,N)$. For any quotient module $N_{0}$ of $N$, we have $\hom_{A}(P,N_{0})\leq\hom_{A}(P,N)$. By Lemma 5.3, we have $\langle{\bf g}_{P},\underline{\dim}N_{0}\rangle=\hom_{A}(N_{0},\tau P)-\hom_{A}(P,N_{0})=-\hom_{A}(P,N_{0})$. Thus
$\langle{\bf g}_{\mathcal{M}},\underline{\dim}N_{0}\rangle$ $=$ $\langle{\bf g}_{M},\underline{\dim}N_{0}\rangle-\langle{\bf g}_{P},\underline{\dim}N_{0}\rangle\quad\;\;(\text{by\;}{\bf g}_{\mathcal{M}}={\bf g}_{M}-{\bf g}_{P})$
$=$ $\langle{\bf g}_{M},\underline{\dim}N_{0}\rangle+\hom_{A}(P,N_{0})$
$\leq$ $\langle{\bf g}_{M},\underline{\dim}N_{0}\rangle+\hom_{A}(P,N)$
[…]
Since $N_{0}$ is an arbitrary quotient module of $N$, we have
$F_{N}[{\bf g}_{\mathcal{M}}]\leq\hom_{A}(N,\tau M)+\hom_{A}(P,N).$
Now let us show the converse inequality. Consider the canonical sequence of $N$ with respect to the torsion pair $(\operatorname{\mathsf{Fac}}M,M^{\bot})$:
(5.13) ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{t}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{f}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0,}$
where $N_{t}\in\operatorname{\mathsf{Fac}}M$ and $N_{f}\in M^{\bot}$.
Since $\mathcal{M}=(M,P)$ is a $\tau$-rigid pair, we have $\operatorname{\mathsf{Hom}}_{A}(P,M)=0$ and thus $\operatorname{\mathsf{Hom}}_{A}(P,N_{t})=0$. Applying the functor $\operatorname{\mathsf{Hom}}_{A}(P,-)$ to (5.13), we see $\operatorname{\mathsf{Hom}}_{A}(P,N)\cong\operatorname{\mathsf{Hom}}_{A}(P,N_{f})$. By Lemma 5.3, we have
$\langle{\bf g}_{P},\underline{\dim}N_{f}\rangle$ $=$ $\hom_{A}(N_{f},\tau P)-\hom_{A}(P,N_{f})$
$=$ $-\hom_{A}(P,N_{f})=-\hom_{A}(P,N).$
By Lemma 5.4, we know that $\langle{\bf g}_{M},\underline{\dim}N_{f}\rangle=\hom_{A}(N,\tau M)$. Thus
$\langle{\bf g}_{\mathcal{M}},\underline{\dim}N_{f}\rangle=\langle{\bf g}_{M},\underline{\dim}N_{f}\rangle-\langle{\bf g}_{P},\underline{\dim}N_{f}\rangle=\hom_{A}(N,\tau M)+\hom_{A}(P,N).$
Since $N_{f}$ is a quotient module of $N$, we have $\langle{\bf g}_{\mathcal{M}},\underline{\dim}N_{f}\rangle\leq F_{N}[{\bf g}_{\mathcal{M}}]$. Thus
$\hom_{A}(N,\tau M)+\hom_{A}(P,N)=\langle{\bf g}_{\mathcal{M}},\underline{\dim}N_{f}\rangle\leq F_{N}[{\bf g}_{\mathcal{M}}].$
Hence, we have $F_{\mathcal{N}}[{\bf g}_{\mathcal{M}}]=F_{N}[{\bf g}_{\mathcal{M}}]=\hom_{A}(N,\tau M)+\hom_{A}(P,N)=E^{\rm proj}(\mathcal{M},\mathcal{N})$. ∎
Remark 5.8.
We remark that the above theorem can be deduced from [fei_2019b, Theorem 3.6]. However, the proofs here and that in [fei_2019b] are quite different.
Corollary 5.9.
Let $\mathcal{M}=\mathcal{M}^{\prime}\oplus\mathcal{M}$$^{\prime\prime}$ be a $\tau$-rigid pair. Then for any decorated $A$-module $\mathcal{N}=\mathcal{N}^{\prime}\oplus\mathcal{N}^{\prime\prime}$, we have
$F_{\mathcal{N}}[{\bf g}_{\mathcal{M}}]=F_{\mathcal{N}}[{\bf g}_{\mathcal{M}^{\prime}}]+F_{\mathcal{N}}[{\bf g}_{\mathcal{M}^{\prime\prime}}]\;\;\;\text{and}\;\;\;F_{\mathcal{N}}[{\bf g}_{\mathcal{M}}]=F_{\mathcal{N}^{\prime}}[{\bf g}_{\mathcal{M}}]+F_{\mathcal{N}^{\prime\prime}}[{\bf g}_{\mathcal{M}}].$
Proof.
It is easy to check that $E^{\rm proj}(\mathcal{M},\mathcal{N})=E^{\rm proj}(\mathcal{M}^{\prime},\mathcal{N})+E^{\rm proj}(\mathcal{M}^{\prime\prime},\mathcal{N})$ and $E^{\rm proj}(\mathcal{M},\mathcal{N})=E^{\rm proj}(\mathcal{M},\mathcal{N}^{\prime})+E^{\rm proj}(\mathcal{M},\mathcal{N}^{\prime\prime})$. Then the desired results follow from Theorem 5.7. ∎
Corollary 5.10.
Let $\mathcal{M}=(M,P)$ and $\mathcal{N}=(N,Q)$ be two $\tau$-rigid pairs in $\operatorname{\mathsf{mod}}\nolimits A$. Then the following statements hold.
  • (i)
    We have $(\mathcal{M}\mid\mid\mathcal{N})_{F}=F_{\mathcal{M}}[{\bf g}_{\mathcal{N}}]+F_{\mathcal{N}}[{\bf g}_{\mathcal{M}}]=E^{\rm sym}(\mathcal{M},\mathcal{N})$.
  • (ii)
    The direct sum $\mathcal{M}\oplus\mathcal{N}$ is a $\tau$-rigid pair in $\operatorname{\mathsf{mod}}\nolimits A$ if and only if $(\mathcal{M}\mid\mid\mathcal{N})_{F}=0$.
Proof.
(i) By Theorem 5.7, we have $F_{\mathcal{N}}[{\bf g}_{\mathcal{M}}]=E^{\rm proj}(\mathcal{M},\mathcal{N})$ and $F_{\mathcal{M}}[{\bf g}_{\mathcal{N}}]=E^{\rm proj}(\mathcal{N},\mathcal{M})$. Thus
$(\mathcal{M}\mid\mid\mathcal{N})_{F}=F_{\mathcal{N}}[{\bf g}_{\mathcal{M}}]+F_{\mathcal{M}}[{\bf g}_{\mathcal{N}}]=E^{\rm proj}(\mathcal{M},\mathcal{N})+E^{\rm proj}(\mathcal{N},\mathcal{M})=E^{\rm sym}(\mathcal{M},\mathcal{N}).$
(ii) Since $E^{\rm sym}(\mathcal{M},\mathcal{N})=\hom_{A}(M,\tau N)+\hom_{A}(Q,M)+\hom_{A}(N,\tau M)+\hom_{A}(P,N)$, we know that $\mathcal{M}\oplus\mathcal{N}$ is $\tau$-rigid if and only if $E^{\rm sym}(\mathcal{M},\mathcal{N})=0$. Then the desirted result follows from (i). ∎
Corollary 5.11.
Let $\mathcal{M}=(M,P)$ be a $\tau$-rigid pair in $\operatorname{\mathsf{mod}}\nolimits A$. Then a module $N\in\operatorname{\mathsf{mod}}\nolimits A$ belongs to $\prescript{\bot}{}{(}\tau M)\cap P^{\bot}$ if and only if $F_{N}[{\bf g}_{\mathcal{M}}]=0$, that is, $\langle{\bf g}_{\mathcal{M}},\underline{\dim}N_{0}\rangle\leq 0$ for any quotient module $N_{0}$ of $N$.
Proof.
By Theorem 5.7, we have $F_{N}[{\bf g}_{\mathcal{M}}]=\hom_{A}(N,\tau M)+\hom_{A}(P,N)$. Then the result follows. ∎
5.2. Newton polytopes of $\tau$-rigid modules
Recall that we use $\mathsf{P}(M)$ to denote the Newton polytope of a module $M\in\operatorname{\mathsf{mod}}\nolimits A$.
Lemma 5.12 ([cao-2025b, Theorem 3.2]).
Let $U$ and $V$ be two $\tau$-rigid modules in $\operatorname{\mathsf{mod}}\nolimits A$. If $\mathsf{P}(U)=\mathsf{P}(V)$, the direct sum $U\oplus V$ remains $\tau$-rigid in $\operatorname{\mathsf{mod}}\nolimits A$.
Proof.
This result is first proved in [cao-2025b]. Here we give a different proof. Since $U$ and $V$ are $\tau$-rigid modules, we know that $E^{\rm proj}(U,U)=\hom_{A}(U,\tau U)=0$ and $E^{\rm proj}(V,V)=\hom_{A}(V,\tau V)=0$. Then by Theorem 5.7, we obtain $F_{U}[{\bf g}_{U}]=E^{\rm proj}(U,U)=0$ and $F_{V}[{\bf g}_{V}]=E^{\rm proj}(V,V)=0$.
Since $U$ and $V$ have the same Newton polytope, we know that $F_{U}[{\bf r}]=F_{V}[{\bf r}]$ for any ${\bf r}\in\mathbb{R}^{n}$. In particular, we have $F_{U}[{\bf g}_{V}]=F_{V}[{\bf g}_{V}]=0$ and $F_{V}[{\bf g}_{U}]=F_{U}[{\bf g}_{U}]=0$. Thus
$(U\mid\mid V)_{F}=F_{U}[{\bf g}_{V}]+F_{V}[{\bf g}_{U}]=0.$
By Corollary 5.10 (ii), we know that $U\oplus V$ is a $\tau$-rigid module. ∎
Lemma 5.13.
Let $\mathcal{M}=\oplus_{i=1}^{n}\mathcal{M}_{i}$ be a basic $\tau$-tilting pair in $\operatorname{\mathsf{mod}}\nolimits A$ and $\mathcal{M}^{\prime}=\mu_{k}(\mathcal{M})=(\oplus_{i\neq k}\mathcal{M}_{i})\oplus\mathcal{M}_{k}^{\prime}$ a left mutation of $\mathcal{M}$. Then we have
  • (i)
    $E^{\rm proj}(\mathcal{M}_{k},\mathcal{M}_{k}^{\prime})=0$ and $E^{\rm proj}(\mathcal{M}_{k}^{\prime},\mathcal{M}_{k})>0$.
  • (ii)
    $F_{\mathcal{M}_{k}^{\prime}}[{\bf g}_{\mathcal{M}_{k}}]=0$ and $F_{\mathcal{M}_{k}}[{\bf g}_{\mathcal{M}_{k}^{\prime}}]>0$.
Proof.
(i) We write $\oplus_{i\neq k}\mathcal{M}_{i}=(U,Q)$, $\mathcal{M}_{k}=(M_{k},Q_{k})$ and $\mathcal{M}_{k}^{\prime}=(M_{k}^{\prime},Q_{k}^{\prime})$. Thus
$\mathcal{M}=(U\oplus M_{k},Q\oplus Q_{k})\;\;\;\text{and}\;\;\;\mathcal{M}^{\prime}=(U\oplus M_{k}^{\prime},Q\oplus Q_{k}^{\prime}).$
Since $\mathcal{M}_{k}$ is indecomposable, we know that either $M_{k}$ or $Q_{k}$ is zero. Similarly, either $M_{k}^{\prime}$ or $Q_{k}^{\prime}$ is zero.
Since the mutation $\mathcal{M}^{\prime}=\mu_{k}(\mathcal{M})$ is a left mutation, we have $\operatorname{\mathsf{Fac}}(U\oplus M_{k}^{\prime})\subsetneq\operatorname{\mathsf{Fac}}(U\oplus M_{k})$. This implies that $M_{k}\neq 0$. Thus $\mathcal{M}_{k}=(M_{k},0)$ and $\mathcal{M}=(U\oplus M_{k},Q)$. Since $\mathcal{M}^{\prime}=(U\oplus M_{k}^{\prime},Q\oplus Q_{k}^{\prime})$ and $\mathcal{M}=(U\oplus M_{k},Q)$ are $\tau$-tilting pairs and by Theorem 3.3 (i), we know that
$\operatorname{\mathsf{Fac}}(U\oplus M_{k}^{\prime})$ $=$ $\prescript{\bot}{}{(\tau U\oplus\tau M_{k}^{\prime})}\cap(Q\oplus Q_{k}^{\prime})^{\bot}=[\prescript{\bot}{}{(\tau U)}\cap Q^{\bot}]\cap[\prescript{\bot}{}{(\tau M_{k}^{\prime})}\cap(Q_{k}^{\prime})^{\bot}],$
$\operatorname{\mathsf{Fac}}(U\oplus M_{k})$ $=$ $\prescript{\bot}{}{(\tau U\oplus\tau M_{k})}\cap Q^{\bot}=[\prescript{\bot}{}{(\tau U)}\cap Q^{\bot}]\cap\prescript{\bot}{}{(\tau M_{k})},$
Since $M_{k}^{\prime}\in\operatorname{\mathsf{Fac}}(U\oplus M_{k}^{\prime})\subsetneq\operatorname{\mathsf{Fac}}(U\oplus M_{k})$ and $M_{k}\in\operatorname{\mathsf{Fac}}(U\oplus M_{k})\setminus\operatorname{\mathsf{Fac}}(U\oplus M_{k}^{\prime})$, we have
$M_{k}^{\prime}\in\prescript{\bot}{}{(\tau M_{k})}\;\;\;\text{and}\;\;\;M_{k}\notin\prescript{\bot}{}{(\tau M_{k}^{\prime})}\cap(Q_{k}^{\prime})^{\bot}.$
Thus we have $E^{\rm proj}(\mathcal{M}_{k},\mathcal{M}_{k}^{\prime})=\hom_{A}(M_{k}^{\prime},\tau M_{k})=0$ and
$E^{\rm proj}(\mathcal{M}_{k}^{\prime},\mathcal{M}_{k})=\hom_{A}(M_{k},\tau M_{k}^{\prime})+\hom_{A}(Q_{k}^{\prime},M_{k})>0.$
(ii) This follows from (i) and Theorem 5.7. ∎
Theorem 5.14.
Let $U$ and $V$ be two $\tau$-rigid modules in $\operatorname{\mathsf{mod}}\nolimits A$. If $U$ and $V$ have the same Newton polytope, i.e., $\mathsf{P}(U)=\mathsf{P}(V)$, then $U\cong V$.
Proof.
By Lemma 5.12, we know that the direct sum $U\oplus V$ is $\tau$-rigid. We will prove that $U\cong V$ by reducing the number $|U\oplus V|$ of iso-classes of indecomposable direct summands of $U\oplus V$. We can assume $|U\oplus V|>0$. Otherwise, we have $U=V=0$ and the desired result holds trivially.
Let $(U\oplus V)^{\flat}$ be the basic $\tau$-rigid module such that $\operatorname{\mathsf{add}}\nolimits(U\oplus V)^{\flat}=\operatorname{\mathsf{add}}\nolimits U\oplus V$. Let $\mathcal{M}=(M,P)$ be the left Bongartz completion of $(U\oplus V)^{\flat}$, which is the basic $\tau$-tilting pair satisfying that $(U\oplus V)^{\flat}$ is a direct summand of $M$ and $\operatorname{\mathsf{Fac}}M=\operatorname{\mathsf{Fac}}(U\oplus V)$.
Let us write $\mathcal{M}=\oplus_{i=1}^{n}\mathcal{M}_{i}$, where each $\mathcal{M}_{i}$ is an indecomposable $\tau$-rigid pair. We can assume that
(5.14) $\mathcal{U}:=(U,0)=\oplus_{i=1}^{n}\mathcal{M}_{i}^{a_{i}}\;\;\;\text{and}\;\;\;\mathcal{V}:=(V,0)=\oplus_{i=1}^{n}\mathcal{M}_{i}^{b_{i}},$
where $a_{i},b_{i}\in\mathbb{Z}_{\geq 0}$.
Since $|U\oplus V|>0$, we know that $\{0\}\subsetneq\operatorname{\mathsf{Fac}}M=\operatorname{\mathsf{Fac}}(U\oplus V)$. Then by Proposition 3.5, there exists a left mutation $\mathcal{M}^{\prime}=(M^{\prime},P^{\prime})=\mu_{k}(\mathcal{M})$ of $\mathcal{M}$ such that $\{0\}\subseteq\operatorname{\mathsf{Fac}}M^{\prime}\subsetneq\operatorname{\mathsf{Fac}}M$.
We claim that $\mathcal{M}_{k}$ is a direct summand of $((U\oplus V)^{\flat},0)$. Otherwise, $((U\oplus V)^{\flat},0)$ is also a direct summand of $\mathcal{M}^{\prime}=(M^{\prime},P^{\prime})$. Then by Remark 3.15 and the fact that $\mathcal{M}=(M,P)$ is the left Bongartz completion of $(U\oplus V)^{\flat}$, we have $\operatorname{\mathsf{Fac}}M\subseteq\operatorname{\mathsf{Fac}}M^{\prime}$. This contradicts $\operatorname{\mathsf{Fac}}M^{\prime}\subsetneq\operatorname{\mathsf{Fac}}M$.
Since $\mathcal{M}_{k}$ is a direct summand of $((U\oplus V)^{\flat},0)$, we can assume that $\mathcal{M}_{k}=(M_{k},0)$, where $M_{k}$ is an indecomposable direct summand of $(U\oplus V)^{\flat}$. Let $\mathcal{M}_{k}^{\prime}$ be the new indecomposable $\tau$-rigid pair obtained in $\mathcal{M}^{\prime}=\mu_{k}(\mathcal{M})$.
By Lemma 5.13, we know that $F_{\mathcal{M}_{k}}[{\bf g}_{\mathcal{M}_{k}^{\prime}}]>0$. For $i\neq k$, we know that $\mathcal{M}_{i}\oplus\mathcal{M}_{k}^{\prime}$ is $\tau$-rigid. Then by Corollary 5.10 (ii), we have $F_{\mathcal{M}_{i}}[{\bf g}_{\mathcal{M}_{k}^{\prime}}]=0$ for any $i\neq k$.
By (5.14) and Proposition 3.8, we have $F_{\mathcal{U}}=\prod_{i=1}^{n}F_{\mathcal{M}_{i}}^{a_{i}}$ and $F_{\mathcal{V}}=\prod_{i=1}^{n}F_{\mathcal{M}_{i}}^{b_{i}}$. Then by Corollary 2.9, we have
$F_{\mathcal{U}}[{\bf g}_{\mathcal{M}_{k}^{\prime}}]$ $=$ $\sum_{i=1}^{n}a_{i}F_{\mathcal{M}_{i}}[{\bf g}_{\mathcal{M}_{k}^{\prime}}]=a_{k}F_{\mathcal{M}_{k}}[{\bf g}_{\mathcal{M}_{k}^{\prime}}],$
$F_{\mathcal{V}}[{\bf g}_{\mathcal{M}_{k}^{\prime}}]$ $=$ $\sum_{i=1}^{n}b_{i}F_{\mathcal{M}_{i}}[{\bf g}_{\mathcal{M}_{k}^{\prime}}]=b_{k}F_{\mathcal{M}_{k}}[{\bf g}_{\mathcal{M}_{k}^{\prime}}].$
Since $U$ and $V$ have the same Newton polytope, we know that $F_{U}[{\bf r}]=F_{V}[{\bf r}]$ for any ${\bf r}\in\mathbb{Z}^{n}$. In particular, we have
$a_{k}F_{\mathcal{M}_{k}}[{\bf g}_{\mathcal{M}_{k}^{\prime}}]=F_{\mathcal{U}}[{\bf g}_{\mathcal{M}_{k}^{\prime}}]=F_{U}[{\bf g}_{\mathcal{M}_{k}^{\prime}}]=F_{V}[{\bf g}_{\mathcal{M}_{k}^{\prime}}]=F_{\mathcal{V}}[{\bf g}_{\mathcal{M}_{k}^{\prime}}]=b_{k}F_{\mathcal{M}_{k}}[{\bf g}_{\mathcal{M}_{k}^{\prime}}].$
Since $F_{\mathcal{M}_{k}}[{\bf g}_{\mathcal{M}_{k}^{\prime}}]>0$, we get $a_{k}=b_{k}$. So we have following decompositions
$U=U^{\prime}\oplus M_{k}^{a_{k}},\quad V=V^{\prime}\oplus M_{k}^{b_{k}}=V^{\prime}\oplus M_{k}^{a_{k}}.$
Note that $U^{\prime}$ and $V^{\prime}$ do not contain $M_{k}$ as a direct summand. In particular, we have
$|U^{\prime}\oplus V^{\prime}|<|U\oplus V|.$
By Corollary 3.10, we know that $\mathsf{P}(U)=\mathsf{P}(U^{\prime})+\mathsf{P}(M_{k}^{a_{k}})$ and $\mathsf{P}(V)=\mathsf{P}(V^{\prime})+\mathsf{P}(M_{k}^{b_{k}})$. Since $\mathsf{P}(U)=\mathsf{P}(V)$ and by Corollary 2.2, we get $\mathsf{P}(U^{\prime})=\mathsf{P}(V^{\prime})$.
In order to show $U\cong V$, it suffices to show $U^{\prime}\cong V^{\prime}$. Now $U^{\prime}$ and $V^{\prime}$ are two $\tau$-rigid modules satisfying $\mathsf{P}(U^{\prime})=\mathsf{P}(V^{\prime})$ and $|U^{\prime}\oplus V^{\prime}|<|U\oplus V|$. We can therefore apply the reduction and complete the proof. ∎
5.3. Newton polytopes of left finite multi-semibricks
A module $M\in\operatorname{\mathsf{mod}}\nolimits A$ is left finite, if the smallest torsion class $\langle M\rangle_{\rm tors}$ containing $M$ is a functorially finite torsion class in $\operatorname{\mathsf{mod}}\nolimits A$. A module $C\in\operatorname{\mathsf{mod}}\nolimits A$ is called a brick, if $\operatorname{\mathsf{End}}_{A}(C)\cong\mathsf{k}$. A set $\Omega$ of iso-classes of bricks is called a semibrick, if $\operatorname{\mathsf{Hom}}_{A}(C,C^{\prime})=0$ for any $C,C^{\prime}\in\Omega$ with $C\neq C^{\prime}$. A semibrick $\Omega$ is left finite, if the smallest torsion class $\langle\Omega\rangle_{\rm tors}$ containing the bricks in $\Omega$ is a functorially finite torsion class.
Proposition 5.15 ([Asai-semibrick, Proposition 2.9]).
The map $\Omega\mapsto\langle\Omega\rangle_{\rm tors}$ gives a bijection from the left finite semibricks in $\operatorname{\mathsf{mod}}\nolimits A$ to functorially finite torsion classes in $\operatorname{\mathsf{mod}}\nolimits A$. In particular, the left finite semibricks in $\operatorname{\mathsf{mod}}\nolimits A$ are in bijection with the basic $\tau$-tilting pairs in $\operatorname{\mathsf{mod}}\nolimits A$.
Let $\mathcal{M}=(M,Q)$ be a basic $\tau$-tilting pair in $\operatorname{\mathsf{mod}}\nolimits A$ and $\Omega$ a left finite semibrick in $\operatorname{\mathsf{mod}}\nolimits A$. The set $\Omega$ is called the labeling semibrick of $\mathcal{M}=(M,Q)$, if $\operatorname{\mathsf{Fac}}M=\langle\Omega\rangle_{\rm tors}$ holds. This name is justified by the following result.
Proposition 5.16 ([Asai-semibrick, Lemma 2.5, Proposition 2.13]).
Let $\mathcal{M}=\oplus_{j=1}^{n}(M_{j},Q_{j})$ a basic $\tau$-tilting pair in $\operatorname{\mathsf{mod}}\nolimits A$. Denote by $I\subseteq[1,n]$ the subset such that $i\in I$ if and only if $\mu_{i}(\mathcal{M})$ is a left mutation of $\mathcal{M}$. Then the following statements hold.
  • (i)
    For each $k\in I$, there exists a unique brick $C_{k}$, called a labeling brick, satisfying that
    $C_{k}\in\operatorname{\mathsf{Fac}}M_{k}\quad\text{and}\quad\operatorname{\mathsf{Hom}}_{A}(M_{j},C_{k})=0\quad\forall j\in[1,n]\setminus k,$
    i.e., $C_{k}$ belongs to $(\oplus_{j\neq k}M_{j})^{\bot}\cap\operatorname{\mathsf{Fac}}M_{k}$.
  • (ii)
    Denote by $\Omega:=\{C_{i}\}_{i\in I}$ the set of labeling bricks for the left mutations of $\mathcal{M}$. Then $\Omega$ is a left finite semibrick and we have $\langle\Omega\rangle_{\rm tors}=\operatorname{\mathsf{Fac}}M$.
Lemma 5.17.
Let $\mathcal{M}=\oplus_{j=1}^{n}\mathcal{M}_{j}$ be a basic $\tau$-tilting pair in $\operatorname{\mathsf{mod}}\nolimits A$ and let $\Omega=\{C_{i}\}_{i\in I}$ be the labeling semibrick of $\mathcal{M}$. Fix a labeling brick $C_{k}\in\Omega$ and let $\mathcal{M}^{\prime}=\mu_{k}(\mathcal{M})=(\oplus_{j\neq k}\mathcal{M}_{j})\oplus\mathcal{M}_{k}^{\prime}$ be the corresponding left mutation of $\mathcal{M}$. Then the following statements hold.
  • (i)
    Denote by $\mathcal{M}_{k}^{\prime}=(M_{k}^{\prime},Q_{k}^{\prime})$. Then $C_{k}\notin\prescript{\bot}{}{(}\tau M_{k}^{\prime})\cap Q_{k}^{\prime\bot}$ and $C_{i}\in\prescript{\bot}{}{(}\tau M_{k}^{\prime})\cap Q_{k}^{\prime\bot}$ for any $i\in I\setminus\{k\}$.
  • (ii)
    We have $F_{C_{k}}[{\bf g}_{{\mathcal{M}}_{k}^{\prime}}]\neq 0$ and $F_{C_{i}}[{\bf g}_{{\mathcal{M}}_{k}^{\prime}}]=0$ for any $i\in I\setminus\{k\}$.
Proof.
(i) Let us write $\mathcal{M}_{j}=(M_{j},Q_{j})$ for $j\in[1,n]$ and $(U,Q):=\oplus_{j\neq k}(M_{j},Q_{j})$. Then
$\mathcal{M}=(U\oplus M_{k},Q\oplus Q_{k})\quad\text{and}\quad\mathcal{M}^{\prime}=(U\oplus M_{k}^{\prime},Q\oplus Q_{k}^{\prime}).$
Since $\mathcal{M}^{\prime}=\mu_{k}(\mathcal{M})$ is a left mutation, we have
$\operatorname{\mathsf{Fac}}(U\oplus M_{k}^{\prime})=\operatorname{\mathsf{Fac}}U\quad\text{and}\quad\operatorname{\mathsf{Fac}}(U\oplus M_{k})=\prescript{\bot}{}{(}\tau U)\cap Q^{\bot}.$
On the other hand, since $\mathcal{M}^{\prime}$ is a $\tau$-tilting pair and by Theorem 3.3 (i), we know that
(5.15) $\operatorname{\mathsf{Fac}}(U\oplus M_{k}^{\prime})=[\prescript{\bot}{}{(}\tau U)\cap Q^{\bot}]\cap[\prescript{\bot}{}{(}\tau M_{k}^{\prime})\cap Q_{k}^{\prime\bot}].$
By Proposition 5.16 (i), we see that $C_{k}\in U^{\bot}\cap\operatorname{\mathsf{Fac}}M_{k}$. On the one hand, we know that
$C_{k}\in\operatorname{\mathsf{Fac}}M_{k}\subseteq\operatorname{\mathsf{Fac}}(U\oplus M_{k})=\prescript{\bot}{}{(}\tau U)\cap Q^{\bot}.$
On the other hand, we know that $C_{k}\in U^{\bot}$ and thus $C_{k}\notin\operatorname{\mathsf{Fac}}U=\operatorname{\mathsf{Fac}}(U\oplus M_{k}^{\prime})$. Then by (5.15), we obtain that $C_{k}\notin\prescript{\bot}{}{(}\tau M_{k}^{\prime})\cap Q_{k}^{\prime\bot}$.
By Proposition 5.16 (i), we know that $C_{i}\in\operatorname{\mathsf{Fac}}M_{i}\subseteq\operatorname{\mathsf{Fac}}U=\operatorname{\mathsf{Fac}}(U\oplus M_{k}^{\prime})$ for $i\in I\setminus\{k\}$. Then by (5.15), we get $C_{i}\in\prescript{\bot}{}{(}\tau M_{k}^{\prime})\cap Q_{k}^{\prime\bot}$.
(ii) This follows from (i) and Corollary 5.11. ∎
Definition 5.18 (Multi-semibrick).
A module $M\in\operatorname{\mathsf{mod}}\nolimits A$ is called a multi-semibrick if $M$ has a decomposition $M\cong\oplus_{i=1}^{r}C_{i}^{a_{i}}$ such that each $C_{i}\in\operatorname{\mathsf{mod}}\nolimits A$ is a brick and $\operatorname{\mathsf{Hom}}_{A}(C_{i},C_{j})=0$ for any $i\neq j$.
For a module $M\in\operatorname{\mathsf{mod}}\nolimits A$, we define its support set $\operatorname{\mathsf{supp}}(M)$ as the set of iso-classes of indecomposable modules which appears as a direct summand of $M$. Clearly, if $M$ is a multi-semibrick, then its support set $\operatorname{\mathsf{supp}}(M)$ is a semibrick.
Lemma 5.19.
Let $U$ and $V$ be two left finite modules in $\operatorname{\mathsf{mod}}\nolimits A$. If $U$ and $V$ have the same Newton polytope, then $\langle U\rangle_{\rm tors}=\langle V\rangle_{\rm tors}$.
Proof.
Since $U$ and $V$ are left finite modules, there are two basic $\tau$-tilting pairs $\mathcal{M}=(M,P)$ and $\mathcal{N}=(N,Q)$ such that $\operatorname{\mathsf{Fac}}M=\langle U\rangle_{\rm tors}$ and $\operatorname{\mathsf{Fac}}N=\langle V\rangle_{\rm tors}$. Since $U\in\operatorname{\mathsf{Fac}}M=\prescript{\bot}{}{(}\tau M)\cap P^{\bot}$, $V\in\operatorname{\mathsf{Fac}}N=\prescript{\bot}{}{(}\tau N)\cap Q^{\bot}$ and by Corollary 5.11, we have $F_{U}[{\bf g}_{\mathcal{M}}]=0$ and $F_{V}[{\bf g}_{\mathcal{N}}]=0$.
Since $U$ and $V$ have the same Newton polytope, we know that $F_{U}[{\bf r}]=F_{V}[{\bf r}]$ for any ${\bf r}\in\mathbb{Z}^{n}$. In particular, we have $F_{U}[{\bf g}_{\mathcal{N}}]=F_{V}[{\bf g}_{\mathcal{N}}]=0$ and $F_{V}[{\bf g}_{\mathcal{M}}]=F_{U}[{\bf g}_{\mathcal{M}}]=0$.
Then by Corollary 5.11 again, we see that $U\in\prescript{\bot}{}{(}\tau N)\cap Q^{\bot}=\operatorname{\mathsf{Fac}}N=\langle V\rangle_{\rm tors}$ and $V\in\prescript{\bot}{}{(}\tau M)\cap P^{\bot}=\operatorname{\mathsf{Fac}}M=\langle U\rangle_{\rm tors}$. Hence, we have $\langle U\rangle_{\rm tors}=\langle V\rangle_{\rm tors}$. ∎
Lemma 5.20.
Let $U$ and $V$ be two left finite multi-semibricks in $\operatorname{\mathsf{mod}}\nolimits A$. If $U$ and $V$ have the same Newton polytope, then $\operatorname{\mathsf{supp}}(U)=\operatorname{\mathsf{supp}}(V)$.
Proof.
Since $U$ and $V$ are two left finite multi-semibricks, we know that $\operatorname{\mathsf{supp}}(U)$ and $\operatorname{\mathsf{supp}}(V)$ are left finite semibricks in $\operatorname{\mathsf{mod}}\nolimits A$. Since $U$ and $V$ have the same Newton polytope and by Lemma 5.19, we know that $\langle U\rangle_{\rm tors}=\langle V\rangle_{\rm tors}$. Since the two left finite semibricks $\operatorname{\mathsf{supp}}(U)$ and $\operatorname{\mathsf{supp}}(V)$ generate the same torsion class $\langle U\rangle_{\rm tors}=\langle V\rangle_{\rm tors}$ and by Proposition 5.15, we have $\operatorname{\mathsf{supp}}(U)=\operatorname{\mathsf{supp}}(V)$. ∎
Theorem 5.21.
Let $U$ and $V$ be two left finite multi-semibricks in $\operatorname{\mathsf{mod}}\nolimits A$. If $U$ and $V$ have the same Newton polytope, then $U\cong V$.
Proof.
By Lemma 5.20, we know that $\operatorname{\mathsf{supp}}(U)=\operatorname{\mathsf{supp}}(V)$. We can assume that $\operatorname{\mathsf{supp}}(U)$ is non-empty. Otherwise, $U=V=0$ and the desired result holds trivially.
Since $U$ and $V$ are left finite and $\operatorname{\mathsf{supp}}(U)=\operatorname{\mathsf{supp}}(V)$, we know that there exists a basic $\tau$-tilting pair $\mathcal{M}=\oplus_{j=1}^{n}\mathcal{M}_{j}=(M,Q)$ such that $\operatorname{\mathsf{Fac}}M=\langle U\rangle_{\rm tors}=\langle V\rangle_{\rm tors}$. Since $\operatorname{\mathsf{supp}}(U)$ is non-empty, we know that $\{0\}\subsetneq\operatorname{\mathsf{Fac}}M$.
Let $I$ be the subset of $[1,n]$ such that $i\in I$ if only if $\mu_{i}(\mathcal{M})$ is a left mutation of $\mathcal{M}$. By Proposition 5.16, we know that the bricks in $\operatorname{\mathsf{supp}}(U)=\operatorname{\mathsf{supp}}(V)$ are indexed by $I$. Say $\operatorname{\mathsf{supp}}(U)=\{C_{i}\}_{i\in I}$. Then we can assume that $U=\oplus_{i\in I}C_{i}^{a_{i}}$ and $V=\oplus_{i\in I}C_{i}^{b_{i}}$, where $a_{i},b_{i}\in\mathbb{Z}_{\geq 1}$.
Let $k\in I$. We know that $\mathcal{M}^{\prime}:=\oplus_{j=1}^{n}\mathcal{M}_{j}^{\prime}=\mu_{k}(\mathcal{M})$ is a left mutation. By Lemma 5.17 (ii), we have
$F_{C_{k}}[{\bf g}_{{\mathcal{M}}_{k}^{\prime}}]\neq 0\quad\text{and}\quad F_{C_{i}}[{\bf g}_{{\mathcal{M}}_{k}^{\prime}}]=0\quad\forall i\in I\setminus\{k\}.$
Since $U=\oplus_{i\in I}C_{i}^{a_{i}}$ and $V=\oplus_{i\in I}C_{i}^{b_{i}}$, we have $F_{U}=\prod_{i\in I}F_{C_{i}}^{a_{i}}$ and $F_{V}=\prod_{i\in I}F_{C_{i}}^{b_{i}}$. Then by Corollary 2.9, we have
$F_{U}[{\bf g}_{{\mathcal{M}}_{k}^{\prime}}]=\sum_{i\in I}a_{i}F_{C_{i}}[{\bf g}_{{\mathcal{M}}_{k}^{\prime}}]=a_{k}F_{C_{k}}[{\bf g}_{{\mathcal{M}}_{k}^{\prime}}]\quad\text{and}\quad F_{V}[{\bf g}_{{\mathcal{M}}_{k}^{\prime}}]$ $=$ $\sum_{i\in I}b_{i}F_{C_{i}}[{\bf g}_{{\mathcal{M}}_{k}^{\prime}}]=b_{k}F_{C_{k}}[{\bf g}_{{\mathcal{M}}_{k}^{\prime}}].$
Since $U$ and $V$ have the same Newton polytope, we have
$a_{k}F_{C_{k}}[{\bf g}_{{\mathcal{M}}_{k}^{\prime}}]=F_{U}[{\bf g}_{{\mathcal{M}}_{k}^{\prime}}]=F_{V}[{\bf g}_{{\mathcal{M}}_{k}^{\prime}}]=b_{k}F_{C_{k}}[{\bf g}_{{\mathcal{M}}_{k}^{\prime}}].$
Since $F_{C_{k}}[{\bf g}_{{\mathcal{M}}_{k}^{\prime}}]\neq 0$, we obtain $a_{k}=b_{k}$. As $k$ is an arbitrary element in $I$, we have $a_{i}=b_{i}$ for any $i\in I$. Thus $U\cong V$. ∎